### v1

# Cross-modal Full-mode Fine-grained Alignment for Text-to-Image Person Retrieval

Abstract.
Text-to-Image Person Retrieval (TIPR) is a cross-modal matching task that aims to retrieve the most relevant person images based on a given text query. The key challenge in TIPR lies in achieving effective alignment between textual and visual modalities within a common latent space. To address this challenge, prior approaches incorporate attention mechanisms for implicit cross-modal local alignment. However, they lack the ability to verify whether all local features are correctly aligned. Moreover, existing methods primarily focus on hard negative samples during model updates, with the goal of refining distinctions between positive and negative pairs, often neglecting incorrectly matched positive pairs. To alleviate these issues, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through explicit fine-grained alignment and existing implicit relational reasoning—hence the term “full-mode”—without requiring additional supervision. Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to rectify unmatched positive sample pairs. A-SDM adaptively pulls the unmatched positive pairs closer in the joint embedding space, thereby achieving more precise global alignment. Additionally, we introduce an Explicit Fine-grained Alignment (EFA) module, which makes up for the lack of verification capability of implicit relational reasoning. EFA strengthens explicit cross-modal fine-grained interactions by sparsifying the similarity matrix and employs a hard coding method for local alignment. Our proposed method is evaluated on three public datasets, achieving state-of-the-art performance among all global matching methods. Our code is available at https://github.com/yinhao1102/FMFA.
Cross-modal retrieval, Person search, Fine-grained alignment
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: tommjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Image searchccs: Computing methodologies Object identification
1. Introduction
Text-to-Image Person Retrieval (TIPR) aims to comprehend natural language descriptions and retrieve the most relevant person image from a large image gallery (Lu et al., 2019). Unlike general image-text retrieval (Sogi et al., 2024; Chen et al., 2024, 2023), which tends to achieve semantic-based matching between text and image, TIPR is specifically designed for identifying individuals. TIPR requires the accurate modeling of fine-grained correspondences between textual and visual modalities, owing to the large intra-class variance and small inter-class difference. This substantial intra-class variation arises from two aspects: (1) visual appearances of the same identity exhibit dramatic variations under different poses, viewpoints, and illumination conditions, and (2) textual descriptions are influenced by differences in phrasing, word order and textual ambiguities. Therefore, the primary challenges in TIPR are how to extract discriminative global representations from image-text pairs and how to achieve precise cross-modal fine-grained alignment. Existing methods for tackling these challenges can be roughly divided into two main categories: global matching methods and local matching methods.
Some global matching methods (Zhang and Lu, 2018; Zheng et al., 2020) obtain discriminative global representations by aligning images and texts, which are projected into a joint embedding space. Their commonly used loss functions include the Cross-Modal Projection Matching (CMPM) loss (Zhang and Lu, 2018) and the Similarity Distribution Matching (SDM) loss (Jiang and Ye, 2023). The CMPM loss focuses on the difference between the scalar projection of image-text pairs and their corresponding matching labels. In comparison, the SDM loss enhances global image-text matching by minimizing the Kullback-Leibler (KL) divergence between the normalized similarity score distribution of image-text pairs and the ground-truth label distribution. In addition, the SDM loss incorporates a temperature hyperparameter to make model updates concentrate on hard negative samples, yet it leads to the neglect of unmatched positive pairs, as shown in Figure 1(a). However, in TIPR, the accurate matching of positive pairs is prioritized over merely distinguishing between positive and negative pairs. Meanwhile, some local matching methods (Bai et al., 2023b; Park et al., 2024; Ergasti et al., 2025) incorporate attention mechanisms to achieve cross-modal fine-grained alignment. For instance, RaSa (Bai et al., 2023b) constructs a cross-modal encoder to generate multimodal representations for subsequent fine-grained alignment. Building on RaSa, MARS (Ergasti et al., 2025) integrates an MAE decoder (He et al., 2022) to reconstruct masked image patch sequences into their original unmasked form, thereby facilitating cross-modal fine-grained alignment. However, these methods rely on attention mechanisms to implicitly aggregate local image-text representations. As a result, they yield only the final multimodal representation, without revealing the details of the aggregation process. Consequently, these implicit aggregation methods make it difficult to determine whether the aggregated multimodal representations correctly encode the corresponding visual and textual information.
(a) Existing global matching methods.
(b) Our adaptive global matching method.
Figure 1. Evolution of global matching methods for text-to-image person retrieval. (a) Existing global matching methods focus on hard negative samples to learn a discriminative boundary in a common latent space, thus enhancing the distinction between positive and negative samples. (b) Our adaptive global matching method builds on a discriminative boundary and concentrates on unmatched positive samples, adaptively pulling them closer to the corresponding query text.
To overcome these issues, we introduce FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through full-mode fine-grained alignment, including explicit fine-grained image-text alignment and existing implicit relational reasoning. Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to ensure the correct matching of positive image-text pairs. Within the joint embedding space, the A-SDM module adaptively pulls positive pairs closer together. In cases of mismatched positive pairs, the A-SDM module adaptively regulates the pulling force based on their relative distance within the joint embedding space, as shown in Figure 1(b), thus improving cross-modal global alignment. Inspired by the observation that multiple image patches usually correspond to one word in the caption (Bica et al., 2024), we introduce an Explicit Fine-Grained Alignment (EFA) module. The EFA module derives multimodal representations through explicit aggregation with a sparse similarity matrix. During this process, the sparse similarity matrix between text and image reflects the contribution of textual and visual representations to the final multimodal representation. To minimize redundancy and reduce the computational cost during training, the EFA module employs hard coding alignment between the aggregated multimodal representation and its original visual and textual representations. These designs enable EFA to achieve cross-modal fine-grained interactions and help the backbone network to extract more discriminative global image-text representations without requiring additional supervision. FMFA is evaluated on three public benchmarks (Li et al., 2017; Ding et al., 2021; Zhu et al., 2021), and it achieves state-of-the-art performance while maintaining high inference speed. Our main contributions are summarized as follows:
  • We propose FMFA to explicitly exploit fine-grained interactions to enhance cross-modal alignment without additional supervision and inference cost.
  • We introduce an adaptive similarity distribution matching module, which is designed to accurately align image-text pairs within a joint embedding space. It adaptively adjusts to narrow the distance between mismatched positive pairs, ensuring more precise matching.
  • We develop an explicit fine-grained alignment module, which leverages the sparse similarity matrix for explicit aggregation and employs a hard coding method in cross-modal fine-grained alignment to minimize redundant information.
2. Related Work
Text-to-Image Person Retrieval (TIPR) was first introduced by Li et al. (Li et al., 2017), who constructed the CUHK-PEDES dataset. The core challenge of TIPR is to learn a common latent space where visual and textual representations are well aligned. Existing methods can be broadly categorized into global and local matching approaches.
Early global methods (Shu et al., 2022; Zheng et al., 2020; Zhu et al., 2021) directly aligned the global representations of images and text in a joint embedding space. Schroff et al. (Schroff et al., 2015) proposed a triplet ranking loss to enforce a margin constraint between positive and negative pairs, and Zhang et al. (Zhang and Lu, 2018) introduced the CMPM/C loss to minimize the discrepancy between the scalar projection of image-text pairs and their labels. However, these global methods lack cross-modal fine-grained interactions, which restrict their ability to capture detailed semantic correspondences. To address this limitation, early local matching methods (Wang et al., 2020; Gao et al., 2021; Shao et al., 2022) explicitly aligned local visual and textual features to achieve fine-grained cross-modal interactions. Nevertheless, they rely on unimodal pre-trained models (e.g., BERT (Devlin et al., 2019) and ResNet (He et al., 2016)), failing to exploit the strong cross-modal alignment capability of recent pre-trained Vision-Language Models (VLMs) (Li et al., 2022; Yan et al., 2023a; Li et al., 2021).
Recent local matching methods (Ergasti et al., 2025; Lu et al., 2025; Qin et al., 2024; Wu et al., 2024; Huang et al., 2025) have benefited greatly from VLMs and introduced VLMs to enhance cross-modal alignment. Park et al. (Park et al., 2024) utilized a modified Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) model as the feature extractor and designed a slot attention-based (Locatello et al., 2020) part discovery module to identify discriminative human parts without extra supervision, while Bai et al. (Bai et al., 2023b) used the align-before-fuse model (Li et al., 2021) as the backbone and introduced a cross-modal encoder for fine-grained alignment. Although effective, these methods involve complex computations during inference, leading to high time and memory costs, which limit their applicability to real-time systems.
On another line of research, several studies (Yang et al., 2023; Shao et al., 2023; Tan et al., 2024) have explored leveraging large-scale image-text pairs in the person Re-IDentification (ReID) domain to VLMs. Zuo et al. (Zuo et al., 2024) utilized CUHK-PEDES and ICFG-PEDES to train an image captioner, aiming to generate comprehensive textual descriptions for pedestrian images. Yang et al. (Yang et al., 2023) employed BLIP-2 (Li et al., 2023) to produce attribute-aware captions for diffusion-generated pedestrian images (Rombach et al., 2022), while Jiang et al. (Jiang et al., 2025) leveraged recent multi-modal large language models (MLLMs), such as Qwen-VL (Bai et al., 2023a) and LLaVA (Liu et al., 2024), to automatically annotate large-scale ReID datasets in a human-like manner. The CLIP models pre-trained on large-scale ReID datasets exhibit strong zero-shot performance. Their compatibility with global matching methods—which relies solely on global features and has a simple inference pipeline—makes them particularly suitable for direct fine-tuning in such settings.
Among the global matching methods, IRRA (Jiang and Ye, 2023) recently attempted to enhance fine-grained interactions by introducing an Implicit Relation Reasoning (IRR) module, utilizing attention mechanisms to capture cross-modal relations. However, this implicit alignment cannot guarantee that all corresponding local features are correctly matched. In light of these limitations, we propose FMFA, which aims to enhance the global matching ability of the model by achieving cross-modal full-mode fine-grained alignment, including explicit fine-grained alignment and implicit relation reasoning.
(a) The architecture of FMFA.
(b) Adaptive Similarity Distribution Matching (A-SDM).
(c) Explicit Fine-grained Alignment (EFA).
Figure 2. Overview of the proposed FMFA framework. (a) Our FMFA consists of a dual-stream feature extraction backbone and four representation learning branches, i.e. Identity Identification (ID loss), Adaptive Similarity Distribution Matching (A-SDM), Explicit Fine-grained Alignment (EFA) and Implicit Relation Reasoning (IRR). The former two are coarse-grained alignment modules and the latter two are fine-grained alignment modules. Modules connected by dashed lines will be removed during inference stage. (b) A-SDM obtains the adaptive weight through the similarity score of unmatched positive pairs, adaptively pulling positive pairs closer and dynamically adjusting the pulling force. (c) EFA achieves fine-grained interaction by hard coding alignment of token or patch embeddings and their corresponding language-grouped vision embedding, which is derived from an aggregated sparse similarity matrix.
3. Method
In this section, we present our proposed FMFA framework. The overview of FMFA is illustrated in Figure 2 and the details are introduced in the following subsections.
3.1. Feature Extraction
Inspired by the success of IRRA (Jiang and Ye, 2023), we use the modified full CLIP (Radford et al., 2021) visual and textual encoders to enhance cross-modal alignment capabilities while reducing inference costs.
Visual Modality. Given an input image $I\in R^{H\times W\times C}$, a CLIP pre-trained ViT model is adopted to obtain the image embeddings. We first split $I$ into a sequence of $N=H\times W/P^{2}$ fixed-sized non-overlapping patches, where $P$ denotes the patch size, and then map the patch sequences to 1D tokens $\left\{f_{i}^{v}\right\}|_{i=1}^{N}$ by a trainable linear projection. With the injection of the positional embedding and the extra [CLS] token, the sequence of tokens $\left\{f_{cls}^{v},f_{1}^{v},...,f_{N}^{v}\right\}$ are input into $L$-layer transformer blocks to model the correlations of each patch. Finally, a linear projection is adopted to map $f_{cls}^{v}$ to the joint image-text embedding space, which serves as the global image representation.
Textual Modality. For an input text $T$, we directly use the CLIP-Xformer textual encoder (Radford et al., 2021) to extract the text embeddings. Following CLIP, the lower-cased byte pair encoding (Sennrich et al., 2016) is firstly employed to tokenize the input text description. The text description is enclosed with [SOS] and [EOS] tokens to indicate the start and end of the sequence. The tokenized text $\{f_{sos}^{t},f_{1}^{t},...,f_{eos}^{t}\}$ is then fed into the textual encoder and exploits the correlations of each patch by masked self-attention. The output of the highest layer of the transformer at the [EOS] token, $f_{eos}^{t}$, is linearly projected into the image-text joint embedding space to obtain the global text representation.
3.2. Adaptive Similarity Distribution Matching
Adopted from IRRA (Jiang and Ye, 2023), we introduce a novel Adaptive Similarity Distribution Matching (A-SDM) module, which aims to adaptively pull the unmatched positive image-text pairs into a joint embedding space, thereby enhancing the cross-modal global matching capability of model.
Given a mini-batch of $B$ image-text pairs, for each global text representation $g_{i}^{t}$ and global image representation $g_{j}^{v}$, we construct a set of image-text representation pairs as $\{(g_{i}^{t},g_{j}^{v}),y_{i,j}\}_{j=1}^{N}$, where $y_{i,j}$ is the matching label. $y_{i,j}=1$ means that $(g_{i}^{t},g_{j}^{v})$ is a matched pair, while $y_{i,j}=0$ means that $(g_{i}^{t},g_{j}^{v})$ is an unmatched pair. Let $cos(\mathbf{u},\mathbf{v})=\mathbf{u}^{\top}\mathbf{v}/\|\mathbf{u}\|\|\mathbf{v}\|$ denotes the cosine similarity of $u$ and $v$. Then, the similarity matrix of matching pairs can be simply calculated by the following softmax function:
(1) $p_{i,j}=\frac{exp(cos(g_{i}^{t},g_{j}^{v})/\tau_{1})}{\sum_{k=1}^{N}exp(cos(g_{i}^{t},g_{k}^{v})/\tau_{1})},$
where $\tau_{1}$ is a temperature hyperparameter which controls the probability distribution peaks. The matching probability $p_{i,j}$ can be viewed as the proportion of the cosine similarity score between $g_{i}^{t}$ and $g_{j}^{v}$ to the sum of cosine similarity scores between $g_{i}^{t}$ and $\{g_{j}^{v}\}_{j=1}^{N}$ in a mini-batch.
Let the ithi-th text $T_{i}$ in the mini-batch be the query text and $I_{i}$ be the corresponding image for $T_{i}$ at rank-$k$, where $k>1$. Different from IRRA (Jiang and Ye, 2023), we propose to calculate adaptive weight based on the similarity between the query text $T_{i}$ and all image representations:
(2) $w_{i}^{t2i}=\alpha\cdot\left[\max_{k}p_{i,k}-p_{i,i}\right]+1,$
where $\alpha$ is a weight factor which reflects the contribution of unmatched image-text pairs to the cross-modal global matching capability of the model. $w_{i}^{t2i}=1$ indicates that the query text $T_{i}$ and its corresponding image $I_{i}$ are successfully matched, while $w_{i}^{t2i}>1$ indicates that $T_{i}$ and $I_{i}$ are unmatched. The A-SDM loss from text to image within a mini-batch is then computed as:
(3) $\mathcal{L}_{t2i}=W^{t2i}*KL(\mathbf{p_{i}\|q_{i}})\\ =\frac{1}{B}\sum_{i=1}^{B}w_{i}^{t2i}\sum_{j=1}^{B}p_{i,j}\log(\frac{p_{i,j}}{q_{i,j}+\epsilon}),$
where $\epsilon$ is a small constant to avoid numerical problems, and $q_{i,j}=y_{i,j}/\sum_{k=1}^{B}$ is the true matching probability.
Symmetrically, the A-SDM loss from image to text $\mathcal{L}_{i2t}$ can be obtained by exchanging the visual and textual modalities. The bi-directional A-SDM loss is calculated by:
(4) $\mathcal{L}_{A-sdm}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}.$
3.3. Explicit Fine-grained Alignment
To fully exploit fine-grained information, it is crucial to bridge the inherent modality differences between visual and textual modalities. Although most existing fine-grained alignment methods based on attention mechanisms have achieved certain results by implicitly aligning local features of images and texts, they cannot directly observe whether the corresponding local features of image and text are effectively aligned. We introduce an explicit cross-modal aggregation method based on the sparse similarity matrix between the image and text local features. To further reduce redundant information and minimize memory and time costs during fine-grained alignment, we use hard coding to align the aggregated language-grouped vision embeddings with both image and text embeddings, as shown in Figure 2(c).
Sparse Similarity Matrix Aggregation. Some methods (Mukhoti et al., 2023; Yao et al., 2022) use loss functions based on fine-grained representations for cross-modal local alignment, which compute the similarity between all image patch embeddings and text token embeddings. Obviously, this approach is computationally and memory-intensive, making it difficult to scale to large batches. Therefore, some methods attempt to sparsify the full-size similarity matrix and only consider the retrieved embeddings that are similar to the query embedding. Most sparse processing methods are based on softmax, but the softmax approach tends to produce a low-entropy distribution from the perspective of gradient flow (Hoffmann et al., 2024). Therefore, we implement sparse similarity matrix aggregation using max-min normalization.
An image $I$ and its corresponding text $T$ are fed into the visual encoder and textual encoder, respectively, as described in Section 3.1. As illustrated in Figure 3(a), the similarity between all image patches and text tokens is obtained by computing the inner product of the last hidden states $\{f_{i}^{t}\}_{i=1}^{L}$ and $\{f_{i}^{v}\}_{i=1}^{N}$ of the text transformer and the vision transformer. $s_{i,j}=f_{i}^{t}\cdot f_{j}^{v}$ represents the similarity between the text token $f_{i}^{t}$ and the image patch $f_{j}^{v}$, where $\cdot$ denotes the inner product. To obtain the aggregation weight, for each token $i$, we first normalize $s_{i,j}$ to [0,1] using the following min-max normalization formula:
(5) $\hat{s}_{i,j}=\frac{s_{i,j}-\min_{k}s_{i,k}}{\max_{k}s_{i,k}-\min_{k}s_{i,k}}.$
We sparsify the normalized similarity matrix to encourage cross-modal interactions between each token and its patches with higher similarity:
(6) $\tilde{s}_{i,j}=\begin{cases}\hat{s}_{i,j}&\mathrm{~if~}\hat{s}_{i,j}\geq\sigma\\ 0&\mathrm{~otherwise}\end{cases},$
where $\sigma$ is the sparsity threshold. $\sigma$ is set to $1/N$, where $N$ is the number of image patches. This ensures that each token has a minimum of one corresponding image patch for alignment. We compute the aggregation weights by:
(7) $agg_{i,j}=\frac{\tilde{s}_{i,j}}{\sum_{m=1}^{M}\tilde{s}_{i,j}},$
where $M$ is the number of image patches retained with high similarity to the token $i$, and $agg_{i,j}$ represents the weight of patch $j$ for aggregating the language-grouped vision embedding corresponding to token $i$. This explicit aggregation strategy ensures a comprehensive interaction between token $i$ and its corresponding patch $j$ during local alignment. In particular, the aggregation weight $agg_{i,j}$ effectively captures the semantic relevance between token $i$ and patch $j$, thereby facilitating precise alignment.
Then, we obtain the corresponding language-grouped vision embedding $e_{i}$ as follows:
(8) ${e}_{i}=\sum_{j=1}^{N}\\ agg_{i,j}\cdot{f}_{j}^{v},$
where $N$ denotes the number of image patches. The resulting set of language-grouped vision embedding ${e}_{i}$ has the same length $L$ as the text token $f_{i}^{t}$.
Hard Coding Alignment. We calculate the similarity between the language-grouped vision embeddings (referred to as joint embeddings) $\{e_{i}\}_{i=1}^{L}$ and their corresponding original text embeddings $\{f_{i}^{t}\}_{i=1}^{L}$ as well as image embeddings $\{f_{i}^{v}\}_{i=1}^{N}$, respectively. To reduce both computational and memory costs, we adopt a hard coding similarity computation between the joint embeddings and their corresponding text and image embeddings. For simplicity, we only present the calculation between the joint embeddings and the text embeddings, while the remaining computations follow a similar and symmetric approach.
(a) Sparse similarity matrix aggregation.
(b) Hard coding alignment.
Figure 3. Illustration of the EFA module. (a) EFA sparifies the similarity matrix between token embeddings and patch embeddings, and obtains the language-grouped vision embeddings by aggregating the sparse similarity with its corresponding patch. (b) EFA obtains hard similarity through hard coding and LSE pooling, where the calculated hard similarity reflects the relationship between the language-grouped vision embeddings and their original token or patch embeddings.
For the text $T$ and its corresponding joint embedding $E$, we calculate the original similarity matrix $O$ between all text tokens $\{f_{i}^{t}\}_{i=1}^{L}$ and their joint embeddings $\{e_{i}\}_{i=1}^{L}$, where $o_{i,j}={f_{i}^{t}}{e_{j}}^{\top}/\|f_{i}^{t}\|\|e_{j}\|$ means the cosine similarity of $f_{i}^{t}$ and $e_{j}$. For the token $f_{i}^{t}$, we compute the weight factor between it and all joint embeddings using the following hard coding way:
(9) $\omega_{i,j}=\begin{cases}1&\mathrm{~if~}j=\underset{j^{\prime}=1\cdots L}{\operatorname*{argmax}}(o_{i,j^{\prime}})\\ 0&otherwise\end{cases}.$
Then, we utilize the LSE pooling (Lee et al., 2018) to compute the hard similarity between text $T$ and its corresponding joint embedding $E$ by:
(10) $\begin{split}hard\_s(T,E)&=LSE-Pooling\left(\sum_{j=1}^{L}\omega_{i,j}o_{i,j}\right)\\ &=\frac{1}{\lambda}\log\sum_{i=1}^{L}\exp\left(\lambda\max_{j=1\cdots L}o_{i,j}\right),\end{split}$
where $\lambda$ is a factor that determines how much to magnify the importance of the most relevant pairs of text embedding and its corresponding joint embedding.
Given a mini-batch of $B$ texts and its corresponding joint embeddings, we obtain the hard coding similarity matrix Hard_SHard\_S according to Eq. (9) and Eq. (10), as illustrated in Figure 3(b). We calculate the EFA loss from the text to its joint embedding, adapted from the triplet ranking loss (Schroff et al., 2015):
(11) $\mathcal{L}_{t2e}=\frac{1}{B}\log\sum_{\mathrm{neg}}\exp\left(\frac{Hard\_S_{\mathrm{neg}}-Hard\_S_{\mathrm{pos}}+margin}{\tau_{2}}\right),$
where $\tau_{2}$ is a temperature hyperparameter that controls the distribution of the loss, and marginmargin is a distance hyperparameter representing the minimum distance difference between positive and negative samples.
Similarly, the EFA loss from the joint embedding to its original text can be computed following Eq. (11), and we can calculate the EFA loss between image and its joint embedding according to Eq. (9), Eq. (10) and Eq. (11). Then, we obtain the full EFA loss by:
(12) $\mathcal{L}_{efa}=\mathcal{L}_{t2e}+\mathcal{L}_{e2t}+\mathcal{L}_{i2e}+\mathcal{L}_{e2i}.$
Theoretical analysis of hard coding: Given visual embeddings of an image $\boldsymbol{V}\in\mathbb{R}^{N\times d}$ and its joint embeddings $\boldsymbol{E}\in\mathbb{R}^{L\times d}$, where $N$ is the number of image patches and $L$ is the length of the joint embeddings, both the hard coding method and the soft coding method calculate the similarity matrix $A$ to obtain the final similarity between $\boldsymbol{V}$ and $\boldsymbol{E}$, which requires the same time complexity $\mathcal{O}(NLd)$. When processing the similarity matrix $A$, our hard coding method calculates the maximum similarity for each row, which has a time complexity of $\mathcal{O}(N)$, as described in Eq. (9) and Eq. (10). However, the soft coding method processes all similarities, resulting in a time complexity of $\mathcal{O}(NL)$. Moreover, the spatial complexity of the hard coding method is $\mathcal{O}(NL)$, which is significantly lower than the spatial complexity of the soft coding method, $\mathcal{O}(Ld)$. This is based on the fact that $N\ll d$. Therefore, our hard coding method can effectively reduce the cost of training time and memory.
3.4. Training Objective
As mentioned, the primary goal of FMFA is to enhance the global and cross-modal local alignment of image-text representations within the joint embedding space. To achieve it, the most commonly used ID loss (Zheng et al., 2020) and IRR loss (Jiang and Ye, 2023), together with the proposed EFA loss and the A-SDM loss, are jointly employed to optimize FMFA. The ID loss directly classifies the global features extracted from each image and text according to their identities, thereby enhancing the global alignment of the model. The IRR loss, based on the Masked Language Modeling (MLM) task (Taylor, 1953), leverages an attention mechanism for implicit cross-modal interaction to obtain a joint embedding, and then predicts the [MASK] text token to enhance the local alignment of the model.
FMFA is trained in an end-to-end manner and the overall training objective is defined as:
(13) $\mathcal{L}=\mathcal{L}_{id}+\mathcal{L}_{irr}+\mathcal{L}_{efa}+\mathcal{L}_{A-sdm}.$
4. Experiments
4.1. Datasets and Settings
In this section, we present the datasets, evaluation metrics and implementation details in our experiments.
Datasets. We use the CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021), and RSTPReid (Zhu et al., 2021) datasets to train and evaluate our FMFA. We follow the data partitions used in IRRA (Jiang and Ye, 2023) to split each dataset into training, validation and test sets, where the ICFG-PEDES dataset only has training and validation sets.
  • CUHK-PEDES is the first benchmark specifically designed for text-to-image person retrieval. It comprises 40,206 images and 80,412 textual descriptions corresponding to 13,003 identities. According to the official split, the training set includes 11,003 identities, with 34,054 images and 68,108 descriptions. The validation and test sets contain 3,078 and 3,074 images, as well as 6,158 and 6,156 textual descriptions, respectively, each covering 1,000 distinct identities.
  • ICFG-PEDES consists of 54,522 images corresponding to 4,102 identities, with each image paired with a single textual description. The dataset is split into a training set and a test set: the training set includes 34,674 image-text pairs from 3,102 identities, while the test set comprises 19,848 pairs covering the remaining 1,000 identities.
  • RSTPReid comprises 20,505 images of 4,101 identities captured by 15 different cameras. Each identity is associated with 5 images taken from different viewpoints, and each image is annotated with 2 textual descriptions. According to the official split, the training, validation, and test sets contain 3,701, 200, and 200 identities, respectively.
Evaluation Metrics. We mainly employ the popular Rank-K metrics (K=1, 5, 10) to measure the retrieval performance. In addition to Rank-K, we also adopt the mean Average Precision (mAP) as an auxiliary retrieval metric to further evaluate performance. Higher evaluation metrics indicate better performance of the model.
[…]
Table 1. The margins utilized in the EFA loss. “T. to E.” means the EFA loss from textual embeddings to the corresponding joint embeddings, and “V. to E.” means the EFA loss from visual embeddings to the corresponding joint embeddings.
CUHK-PEDES ICFG-PEDES RSTPReid
T. to E. 0.1 0.2 0.2
V. to E. 0.1 1.0 0.8
4.2. Comparison with State-of-the-Art Methods
In this section, we present a comparison with state-of-the-art methods (e.g., NAM (Tan et al., 2024) and HAM (Jiang et al., 2025)) on three public benchmark datasets. Based on the backbone type, we categorize the baselines into two groups, as shown in Table 2, Table 3, and Table 4: those employing VL-Backbones w/o ReID-domain pre-training and with ReID-domain pre-training. Furthermore, according to whether local features are utilized during inference, the baselines are further divided into local matching methods and global matching methods (denoted as “L” and “G” in the “Type” column, respectively). Note that the baseline model in Table 2, Table 3 and Table 4 is denoted as IRRAR, which represents the performance of our reimplementation of the IRRA model. CLIP refers to the CLIP-ViT-B/16 model fine-tuned with the InfoNCE loss (Oord et al., 2018).
Table 2. Performance comparisons with state-of-the-art methods on the CUHK-PEDES dataset. “G”, “L” and “P” in the “Type” column stand for global-matching method, local-matching method and pre-trained model with ReID-domain respectively. “Image Enc.” and “Text Enc.” mean the backbone of image encoder and text encoder respectively. “IRRAR” means the model that we reproduce.
Type Method Ref. Image Enc. Text Enc. Rank-1 Rank-5 Rank-10 mAP
VL-Backbones w/o ReID-domain pre-training:
G LGUR (Shao et al., 2022) MM22 ResNet50 BERT 65.25 83.12 89.00 -
IVT (Shu et al., 2022) ECCV22 ViT-B/16 BERT 65.59 83.11 89.21 -
CLIP (Radford et al., 2021) ICML21 CLIP-ViT CLIP-Xformer 68.19 86.47 91.47 61.12
DM-Adapter (Liu et al., 2025) AAAI25 CLIP-ViT CLIP-Xformer 72.17 88.74 92.85 64.33
IRRAR (Jiang and Ye, 2023) CVPR23 CLIP-ViT CLIP-Xformer 73.45 89.38 93.69 66.13
TBPS-CLIP (Cao et al., 2024) AAAI24 CLIP-ViT CLIP-Xformer 73.54 88.19 92.35 65.38
FMFA (ours) CLIP-ViT CLIP-Xformer 74.16 90.12 94.10 66.66
L ACSA (Ji et al., 2022) TMM22 Swin-B BERT 63.56 81.49 87.70 -
Han et al. (Han et al., 2021) arXiv21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08
PLOT (Park et al., 2024) ECCV24 CLIP-ViT CLIP-Xformer 75.28 90.42 94.12 -
RaSa (Bai et al., 2023b) IJCAI23 CLIP-ViT BERT-base 76.51 90.29 94.25 69.38
PTMI (Lu et al., 2025) TIFS25 CLIP-ViT CLIP-Xformer 76.02 89.93 94.14 70.85
APTM (Yang et al., 2023) MM23 Swin-B BERT-base 76.53 90.04 94.15 66.91
SCVD (Wei et al., 2024) TCSVT24 CLIP-RN50 CLIP-Xformer 76.72 90.38 94.89 -
VL-Backbones with ReID-domain pre-training:
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) ICCV23 CLIP-ViT CLIP-Xformer 74.37 89.51 93.97 66.60
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) NeurIPS24 CLIP-ViT CLIP-Xformer 74.25 89.49 93.68 66.52
NAM (Tan et al., 2024) + IRRAR CVPR24 CLIP-ViT CLIP-Xformer 76.67 91.11 94.60 68.42
NAM (Tan et al., 2024) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.23 91.33 94.75 68.53
HAM (Jiang et al., 2025) + IRRAR CVPR25 CLIP-ViT CLIP-Xformer 77.32 91.20 94.95 68.87
HAM (Jiang et al., 2025) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.46 91.36 95.01 68.89
Performance Comparisons on CUHK-PEDES. We evaluate FMFA on the widely used CUHK-PEDES benchmark, as shown in Table 2. When using the VL-Backbones without ReID-domain pre-training, FMFA outperforms all state-of-the-art global matching methods with 74.16% Rank-1 accuracy and 66.66% mAP, and further exceeds IRRA by 0.74% and 0.41% in Rank-5 and Rank-10, respectively. When adopting the VL-Backbones with ReID-domain pre-training, FMFA maintains its superiority, and achieves Rank-5 accuracy exceeding 95% with the HAM-based backbone. Notably, FMFA with NAM-based backbone attains 91.33% in Rank-5, outperforming IRRA with the HAM-based backbone by 0.13%.
Table 3. Performance comparisons with state-of-the-art methods on the RSTPReid dataset.
Type Method Rank-1 Rank-5 Rank-10 mAP
VL-Backbones w/o ReID-domain pre-training:
G DSSL (Zhu et al., 2021) 39.05 62.60 73.95 -
IVT (Shu et al., 2022) 46.70 70.00 78.80 -
CLIP (Radford et al., 2021) 54.05 80.70 88.00 43.41
IRRAR (Jiang and Ye, 2023) 59.50 81.80 88.85 47.44
DM-Adapter (Liu et al., 2025) 60.00 82.10 87.90 47.37
FMFA (ours) 61.05 83.85 89.80 48.22
L ACSA (Lu et al., 2025) 48.40 71.85 81.45 -
CFine (Yan et al., 2023a) 50.55 72.50 81.60 -
PLOT (Park et al., 2024) 61.80 82.85 89.45 -
RaSa (Bai et al., 2023b) 66.90 86.50 91.35 52.31
APTM (Yang et al., 2023) 67.50 85.70 91.45 52.56
VL-Backbones with ReID-domain pre-training:
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 62.20 83.30 89.75 48.33
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 64.35 83.75 91.00 50.93
NAM (Tan et al., 2024) + IRRAR 68.25 86.75 92.30 52.92
NAM (Tan et al., 2024) + FMFA (ours) 68.70 87.05 92.35 53.14
HAM (Jiang et al., 2025) + IRRAR 71.35 87.60 93.05 55.40
HAM (Jiang et al., 2025) + FMFA (ours) 71.80 88.05 93.15 55.72
Performance Comparisons on RSTPReid. We evaluate FMFA on the newest benchmark, RSTPReid, as shown in Table 3. Using the VL-Backbones without ReID-domain pre-training, FMFA achieves competitive performance, attaining 61.05% Rank-1, 83.85% Rank-5, 89.80% Rank-10, and 48.22% mAP, respectively, outperforming IRRA by 1.55% in Rank-1 and 2.05% in Rank-5. When adopting the VL-Backbones with ReID-domain pre-training, our method achieves further gains, exceeding IRRA by 0.45% in Rank-1 with both the NAM-based and HAM-based backbones. Notably, FMFA achieves Rank-5 accuracy higher than 88% with the HAM-based backbone.
Table 4. Performance comparisons with state-of-the-art methods on the ICFG-PEDES dataset.
Type Method Rank-1 Rank-5 Rank-10 mAP
VL-Backbones w/o ReID-domain pre-training:
G Dual Path (Zheng et al., 2020) 38.99 59.44 68.41 -
IVT (Shu et al., 2022) 56.04 73.60 80.22 -
CLIP (Radford et al., 2021) 56.74 75.72 82.26 31.84
DM-Adapter (Liu et al., 2025) 62.64 79.53 85.32 36.50
IRRAR (Jiang and Ye, 2023) 63.48 80.16 85.78 38.20
FMFA (ours) 64.29 80.48 85.93 39.43
L SSAN (Ding et al., 2021) 54.23 72.63 79.53 -
ISANet (Yan et al., 2023b) 57.73 75.42 81.72 -
CFine (Yan et al., 2023a) 60.83 76.55 82.42 -
RaSa (Bai et al., 2023b) 65.28 80.40 85.12 41.29
PLOT (Park et al., 2024) 65.76 81.39 86.73
VL-Backbones with ReID-domain pre-training:
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 64.50 80.24 85.74 38.22
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 65.79 81.94 87.32 39.43
NAM (Tan et al., 2024) + IRRAR 66.34 81.94 86.73 40.14
NAM (Tan et al., 2024) + FMFA (ours) 66.58 81.94 87.04 40.17
HAM (Jiang et al., 2025) + IRRAR 68.21 83.28 88.04 41.72
HAM (Jiang et al., 2025) + FMFA (ours) 68.37 83.28 88.10 41.76
Performance Comparisons on ICFG-PEDES. We also evaluate FMFA on another benchmark, ICFG-PEDES, as shown in Table 4. Using VL-Backbones without ReID-domain pre-training, FMFA achieves the best performance across all evaluation metrics, attaining 64.29% Rank-1 and 39.43% mAP. Compared with IRRA, FMFA shows a notable improvement of 0.81% in Rank-1 and 1.23% in mAP, which is meaningful for practical applications. When adopting VL-Backbones with ReID-domain pre-training, FMFA yields slight gains, outperforming IRRA by 0.24% and 0.16% in Rank-1 with the NAM-based and HAM-based backbones, respectively.
In summary, our FMFA achieves the best performance across all evaluation metrics on three common public benchmarks. To the best of our knowledge, FMFA is the best method for all global matching methods. This demonstrates the generalizability and robustness of our proposed method.
Table 5. Ablation study on each component of FMFA on CUHK-PEDES, ICFG-PEDES and RSTPReid.
No. Methods Components CUHK-PEDES ICFG-PEDES RSTPReid
SDM A-SDM EFA Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10
0 Baseline $\checkmark$ 73.45 89.38 93.69 63.48 80.16 85.78 59.50 81.80 88.85
1 +A-SDM $\checkmark$ 74.04 89.86 93.89 64.26 80.59 85.90 60.25 81.45 88.70
2 +EFA $\checkmark$ $\checkmark$ 73.73 89.64 94.04 63.77 80.39 85.86 60.45 82.30 89.25
3 FMFA $\checkmark$ $\checkmark$ 74.16 90.12 94.10 64.29 80.48 85.93 61.05 83.85 89.80
4.3. Ablation Study
In this section, we analyze our proposed components in the FMFA framework. For simplicity, we omit the components of $\mathcal{L}_{id}$ and the IRR module that were proposed by IRRA and used in all experiments. Only one of SDM and A-SDM can be used at the same time.
To fully demonstrate the impact of our proposed components in FMFA, we conduct an empirical analysis on three common datasets (i.e., CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021) and RSTPReid (Zhu et al., 2021)). The Rank-1, Rank-5 and Rank-10 accuracies(%) are shown in Table 5.
Effect of The A-SDM Module. To evaluate the contribution of the Adaptive Similarity Distribution Matching (A-SDM) module, we perform ablation experiments by replacing the A-SDM module with the SDM module, keeping all hyperparameters unchanged. Specifically, as shown in Table 5, replacing A-SDM with SDM leads to a decrease in the Rank-1 accuracy by 0.59%, 0.78%, and 0.75% across the three datasets, as observed in No. 0 vs. No. 1. Additionally, all evaluation metrics on the CUHK-PEDES and ICFG-PEDES datasets degrade, further confirming the superiority of A-SDM. Moreover, when combined with the EFA module, the advantage of A-SDM becomes even more pronounced. As shown in No. 2 vs. No. 3, replacing the A-SDM module with the SDM module results in 1.26% and 1.00% decrease in Rank-1 and Rank-5 on the CUHK-PEDES dataset, respectively, and a 1.55% drop in the Rank-5 accuracy on the RSTPReid dataset. These results collectively validate the consistent and significant contribution of the A-SDM module to overall performance.
Effect of The EFA Module. To enhance the global matching capability of model, the Explicit Fine-grained Alignment (EFA) module introduces fine-grained cross-modal interaction based on a sparse similarity matrix. The effectiveness of the EFA module is demonstrated through the experimental comparisons between No. 0 vs. No. 2 and No. 1 vs. No. 3. Specifically, as shown in No. 1 vs. No. 3, removing the EFA module from FMFA leads to a performance drop of 0.26% and 0.21% in Rank-5 and Rank-10 on the CUHK-PEDES dataset, and a more significant decline of 2.40% and 1.10% in Rank-5 and Rank-10 on the RSTPReid dataset. However, EFA causes a 0.11% drop in Rank-5 on ICFG-PEDES, suggesting that its sparse and hard coding strategy, which focuses only on the most relevant patches, may overlook other informative ones and lead to information loss. Notably, this comparison also reflects the joint ablation of EFA and the A-SDM module, further verifying the complementary effect between the two modules. Moreover, to further validate the individual contribution of the EFA module, we design an additional experiment that retains the SDM module while removing only the EFA module, as observed in No. 0 vs. No. 2. In this setting, the absence of EFA results in 0.38%, 0.29%, and 0.95% drops in Rank-1 on CUHK-PEDES, ICFG-PEDES and RSTPReid, respectively. These comparisons demonstrate the effectiveness of the EFA module.
(a) The weight factor $\alpha$ in Eq. (2).
(b) The sparsity threshold $\sigma$ in Eq. (6), where $N$ is the number of image patches.
(c) The LSE factor $\lambda$ in Eq. (10).
Figure 4. The sensitivity analysis of hyperparameters of FMFA on the CUHK-PEDES dataset.
4.4. Parameter Study
We conduct the parameter analysis on the CUHK-PEDES dataset, including three hyperparameters (i.e., the weight factor $\alpha$, the sparsity threshold $\sigma$ and the factor $\lambda$) and the weight of the proposed loss functions (i.e., $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$). While we analyze the special parameter, other parameters are kept the same as those described in Section 4.1.
Hyperparameters Analysis. As shown in Figure 4(a), we vary the weight factor $\alpha$ from 1 to 20. The results on the CUHK-PEDES dataset indicate that setting $\alpha$ to 10 yields the best performance across all metrics, suggesting that the adaptive pull force on unmatched positive pairs is optimal. However, when $\alpha$ is increased to 20, the performance drops significantly because the pull force becomes excessively strong, causing unmatched positive pairs to fail to align properly and instead over-pulling mismatched positives, which leads to false positives. We further vary the sparsity threshold $\sigma$ from $0.1/N$ to $20/N$ and the LSE factor $\lambda$ from 0.1 to 10, where $N$ denotes the number of image patches. As illustrated in Figure 4(b) and Figure 4(c), setting $\sigma$ to $1/N$ and $\lambda$ to 1 yields the best overall performance.
Loss Function Weights. We conduct a series of experiments on the CUHK-PEDES dataset to analyze the effect of the weights of the two proposed loss functions, $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$, varying them from 0.1 to 10.0, as shown in Table 6. The results show that setting both weights to 1 yields the best performance. Notably, increasing the weight of $\mathcal{L}_{efa}$ to 5.0 or 10.0 results in gradient explosion, making the model untrainable and causing all metrics to drop below 1.
Table 6. The sensitivity analysis of the weights of $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$ on the CUHK-PEDES dataset.
Weight $\mathcal{L}_{A-sdm}$ $\mathcal{L}_{efa}$
Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10
0.1 71.47 88.62 93.03 68.32 85.75 91.04
0.5 71.23 88.06 93.16 58.49 80.05 87.13
1.0 74.16 90.12 94.10 74.16 90.12 94.10
5.0 71.86 88.27 93.12 0.09 0.19 0.32
10.0 72.45 88.62 93.24 0.09 0.19 0.32
Figure 5. Visualization of similarity maps before and after the sparse process in the EFA module.
4.5. Qualitative Results
Visualization of The Sparse Process. We visualize the similarity maps before and after the sparse process (i.e., Eq. (5) and Eq. (6)) in the EFA module, as shown in Figure 5. Patches with high similarity (greater than 0.75) are preserved, whereas those with low similarity (below 0.5) are suppressed to values under 0.25 and thus omitted during aggregation (left in Figure 5). When most patches in an image exhibit relatively high similarity (around 0.75), the sparse process retains only the patches with the highest similarity, while pushing the similarity of the remaining patches below 0.5 (right in Figure 5). This ensures that the subsequent aggregation focuses only on the most relevant patches, reducing computational and memory costs while maintaining performance. Unlike implicit aggregation methods based on attention mechanisms, EFA explicitly aggregates image patches and text tokens, allowing us to observe whether the most relevant patches are effectively aggregated by visualizing the similarity between patches and tokens.
Inference Time Comparison We compare the inference time between FMFA and recent local matching methods (e.g., PLOT (Park et al., 2024) and RaSa (Bai et al., 2023b)) on the test sets of three datasets, as shown in Table 7. As a global matching method, FMFA only computes global features during inference, thus achieving a higher inference speed than local matching methods. Moreover, the efficiency advantage of FMFA becomes more pronounced as the size of the test set increases. For instance, the inference time of FMFA vs. PLOT increases from 3s vs. 5s on RSTPReid to 50s vs. 91s on ICFG-PEDES.
Table 7. Comparison of the inference time (s) between FMFA and recent local matching methods.
Method RSTPReid CUHK-PEDES ICFG-PEDES
FMFA (ours) 3 7 50
PLOT (Park et al., 2024) 5 16 91
RaSa (Bai et al., 2023b) 388 1168 3871
(a) A man wearing a gray and black shirt, a pair of black pants and a bag over his right shoulder.
(b) The man has short dark hair and is wearing a white t-shirt, dark pants and a black backpack.
(c) The woman has a long pony tail and is wearing a teal color shirt and matching shorts or skirt.
(d) Female holding her phone with long black hair, wearing jogging clothes and tennis shoes carrying bags.
Figure 6. Comparison of top-5 retrieved results on CUHK-PEDES between baseline and FMFA for each text query. The image corresponding to query text, along with the matched and mismatched images, are highlighted with black, green and red rectangles, respectively.
Visualization of Top-5 Retrieval Results. Figure 6 compares the top-5 retrieval results between the baseline IRRAR and our proposed FMFA on the CUHK-PEDES dataset. As shown in Figure 6, FMFA achieves more accurate retrieval results, successfully retrieving images that the baseline fails to match. For query texts where the baseline performs well, FMFA further improves performance by retrieving more relevant pedestrian images (e.g., Figure 6(a) and Figure 6(c)). Even for hard negative samples, where the baseline struggles to retrieve the correct image, FMFA still enhances the similarity between positive pairs (e.g., Figure 6(b) and Figure 6(d)). This is because our proposed FMFA focuses on the unmatched positive pairs and adaptively pulls the positive pairs closer. More comparisons of the top-5 retrieved results between the baseline and FMFA on the ICFG-PEDES and RSTPReid datasets are provided in Figure 7 and Figure 8.
(a) A man in his forties with short black hair is wearing a mid-length grey trench coat over a light-colored collared shirt. He is also wearing a regular-fit black pants.
(b) A woman with black hair tied at the back is wearing a black knee-length overcoat and black pants. She is wearing black shoes and holding two white shopping bags in her hand.
(c) A woman with medium length black hair is wearing a black jacket with black pants and black sneakers. She’s also carrying a white shopping bag.
(d) A middle-aged man with black medium length hair with receding hairline and he is wearing a green hooded jacket with black pants paired with black shoes.
Figure 7. Comparison of top-5 retrieved results on ICFG-PEDES between baseline and FMFA for each text query. The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
(a) A person is wearing a short dark jacket, a pair of green overalls and a pair of sneakers with blue edging.
(b) The man who wears the red and black overcoat is wearing a pair of blue jeans and carrying a black backpack.
(c) The man is wearing a gray coat with the hood. He wears a pair of black trousers and a pair of black shoes.And he is carrying a dark single shoulder bag.
(d) A man walking on the street is wearing a navy blue jacket,loose jeans and a pair of sneakers. He puts his jacket’s hat on the head.And a bag and some paper in his hand.
Figure 8. Comparison of top-5 retrieved results on RSTPReid between baseline and FMFA for each text query. The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
5. Conclusion
In this paper, we propose a cross-modal Full-Mode Fine-grained Alignment (FMFA) framework to learn discriminative global text-image representations through full-mode fine-grained alignment, including explicit fine-grained alignment and existing implicit relational reasoning. We design an Adaptive Similarity Distribution Matching (A-SDM) module to concentrate on unmatched positive pairs, adaptively pulling them closer. In addition, to achieve cross-modal fine-grained alignment, we introduce an Explicit Fine-grained Alignment (EFA) module, which explicitly aggregates local text and image representations based on the sparse similarity matrix and employs a hard coding method. These modules work collaboratively to align images and text into a joint embedding space. Extensive experiments on three datasets demonstrate the superiority and effectiveness of our proposed FMFA framework.
Limitation. The fixed threshold in the sparse process only keeps the most relevant patches, which may result in the loss of semantic information and limit the effective aggregation of local features, thereby affecting the overall performance of our model. Incorporating adaptive methods that capture complete semantic information (e.g., tree transformer (Wang et al., 2019)) could further enhance our model.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62302080), National Foreign Expert Project of China (No. S20240327), Sichuan Science and Technology Program (No. 2025HJRC0021) and Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No. BX202312).

### v2

@@
 Abstract.
-Text-to-Image Person Retrieval (TIPR) is a cross-modal matching task that aims to retrieve the most relevant person images based on a given text query.
-The key challenge in TIPR lies in achieving effective alignment between textual and visual modalities within a common latent space.
+Text-to-Image Person Retrieval (TIPR) is a cross-modal matching task designed to identify the person images that best correspond to a given textual description.
+The key difficulty in TIPR is to realize robust correspondence between the textual and visual modalities within a unified latent representation space.
 To address this challenge, prior approaches incorporate attention mechanisms for implicit cross-modal local alignment.
 However, they lack the ability to verify whether all local features are correctly aligned.
-Moreover, existing methods primarily focus on hard negative samples during model updates, with the goal of refining distinctions between positive and negative pairs, often neglecting incorrectly matched positive pairs.
-To alleviate these issues, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through explicit fine-grained alignment and existing implicit relational reasoning—hence the term “full-mode”—without requiring additional supervision.
-Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to rectify unmatched positive sample pairs.
+Moreover, existing methods tend to emphasize the utilization of hard negative samples during model optimization to strengthen discrimination between positive and negative pairs, often neglecting incorrectly matched positive pairs.
+To mitigate these problems, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through explicit fine-grained alignment and existing implicit relational reasoning—hence the term “full-mode”—without introducing extra supervisory signals.
+In particular, we propose an Adaptive Similarity Distribution Matching (A-SDM) module to rectify unmatched positive sample pairs.
 A-SDM adaptively pulls the unmatched positive pairs closer in the joint embedding space, thereby achieving more precise global alignment.
@@
 EFA strengthens explicit cross-modal fine-grained interactions by sparsifying the similarity matrix and employs a hard coding method for local alignment.
-Our proposed method is evaluated on three public datasets, achieving state-of-the-art performance among all global matching methods.
-Our code is available at https://github.com/yinhao1102/FMFA.
+We evaluate our method on three public datasets, where it attains state-of-the-art results among all global matching methods.
+The code for our method is publicly accessible at https://github.com/yinhao1102/FMFA.
 Cross-modal retrieval, Person search, Fine-grained alignment
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: tommjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Image searchccs: Computing methodologies Object identification
+copyright: acmlicensedjournal: TOMMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3786798ccs: Information systems Image searchccs: Computing methodologies Object identification
 1. Introduction
-Text-to-Image Person Retrieval (TIPR) aims to comprehend natural language descriptions and retrieve the most relevant person image from a large image gallery (Lu et al., 2019).
-Unlike general image-text retrieval (Sogi et al., 2024; Chen et al., 2024, 2023), which tends to achieve semantic-based matching between text and image, TIPR is specifically designed for identifying individuals.
+Text-to-Image Person Retrieval (TIPR) seeks to understand natural language descriptions and identify the most relevant person image within a large gallery (lu2019vilbert).
+Unlike general image-text retrieval (wang2017adversarial; sogi2024object; chen2024make; chen2023vilem; wang2025geometric), which tends to achieve semantic-based matching between text and image, TIPR is specifically designed for identifying individuals.
 TIPR requires the accurate modeling of fine-grained correspondences between textual and visual modalities, owing to the large intra-class variance and small inter-class difference.
@@
 Existing methods for tackling these challenges can be roughly divided into two main categories: global matching methods and local matching methods.
-Some global matching methods (Zhang and Lu, 2018; Zheng et al., 2020) obtain discriminative global representations by aligning images and texts, which are projected into a joint embedding space.
-Their commonly used loss functions include the Cross-Modal Projection Matching (CMPM) loss (Zhang and Lu, 2018) and the Similarity Distribution Matching (SDM) loss (Jiang and Ye, 2023).
-The CMPM loss focuses on the difference between the scalar projection of image-text pairs and their corresponding matching labels.
-In comparison, the SDM loss enhances global image-text matching by minimizing the Kullback-Leibler (KL) divergence between the normalized similarity score distribution of image-text pairs and the ground-truth label distribution.
+Some global matching methods (zhang2018deep; zheng2020dual) obtain discriminative global representations by aligning images and texts, which are projected into a joint embedding space.
+Their widely adopted loss functions include the Cross-Modal Projection Matching (CMPM) loss (zhang2018deep) and the Similarity Distribution Matching (SDM) loss (jiang2023cross).
+The CMPM loss highlights the gap between the scalar projections of image-text pairs and their matched label indicators.
+In comparison, the SDM loss boosts global matching performance by minimizing the Kullback-Leibler (KL) divergence between the normalized similarity profile of image-text pairs and the true label distribution.
 In addition, the SDM loss incorporates a temperature hyperparameter to make model updates concentrate on hard negative samples, yet it leads to the neglect of unmatched positive pairs, as shown in Figure 1(a).
 However, in TIPR, the accurate matching of positive pairs is prioritized over merely distinguishing between positive and negative pairs.
-Meanwhile, some local matching methods (Bai et al., 2023b; Park et al., 2024; Ergasti et al., 2025) incorporate attention mechanisms to achieve cross-modal fine-grained alignment.
-For instance, RaSa (Bai et al., 2023b) constructs a cross-modal encoder to generate multimodal representations for subsequent fine-grained alignment.
-Building on RaSa, MARS (Ergasti et al., 2025) integrates an MAE decoder (He et al., 2022) to reconstruct masked image patch sequences into their original unmasked form, thereby facilitating cross-modal fine-grained alignment.
+Meanwhile, some local matching methods (bai2023rasa; park2024plot; ergasti2025mars) incorporate attention mechanisms to achieve cross-modal fine-grained alignment.
+For instance, RaSa (bai2023rasa) constructs a cross-modal encoder to generate multimodal representations for subsequent fine-grained alignment.
+Building on RaSa, MARS (ergasti2025mars) integrates a Masked AutoEncoder (MAE) decoder (he2022masked) to reconstruct masked image patch sequences into their original unmasked form, thereby facilitating cross-modal fine-grained alignment.
 However, these methods rely on attention mechanisms to implicitly aggregate local image-text representations.
@@
 (b) Our adaptive global matching method builds on a discriminative boundary and concentrates on unmatched positive samples, adaptively pulling them closer to the corresponding query text.
-To overcome these issues, we introduce FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through full-mode fine-grained alignment, including explicit fine-grained image-text alignment and existing implicit relational reasoning.
+To remedy these concerns, we propose FMFA, a cross-modal Full-Mode Fine-grained Alignment framework, which enhances global matching through full-mode fine-grained alignment, including explicit fine-grained image-text alignment and existing implicit relational reasoning.
 Specifically, we design an Adaptive Similarity Distribution Matching (A-SDM) module to ensure the correct matching of positive image-text pairs.
@@
 In cases of mismatched positive pairs, the A-SDM module adaptively regulates the pulling force based on their relative distance within the joint embedding space, as shown in Figure 1(b), thus improving cross-modal global alignment.
-Inspired by the observation that multiple image patches usually correspond to one word in the caption (Bica et al., 2024), we introduce an Explicit Fine-Grained Alignment (EFA) module.
+Based on the insight that each word in a caption can be associated with several image patches (bica2024improving), we introduce an Explicit Fine-Grained Alignment (EFA) module.
 The EFA module derives multimodal representations through explicit aggregation with a sparse similarity matrix.
@@
 To minimize redundancy and reduce the computational cost during training, the EFA module employs hard coding alignment between the aggregated multimodal representation and its original visual and textual representations.
-These designs enable EFA to achieve cross-modal fine-grained interactions and help the backbone network to extract more discriminative global image-text representations without requiring additional supervision.
-FMFA is evaluated on three public benchmarks (Li et al., 2017; Ding et al., 2021; Zhu et al., 2021), and it achieves state-of-the-art performance while maintaining high inference speed.
-Our main contributions are summarized as follows:
-  • We propose FMFA to explicitly exploit fine-grained interactions to enhance cross-modal alignment without additional supervision and inference cost.
-  • We introduce an adaptive similarity distribution matching module, which is designed to accurately align image-text pairs within a joint embedding space.
+These designs allow EFA to realize fine-grained cross-modal interactions and assist the backbone network in learning more distinctive global image-text representations without introducing additional supervision.
+FMFA is evaluated on three public benchmarks (li2017person; ding2021semantically; zhu2021dssl), and it attains competitive top-level performance along with high inference efficiency.
+We highlight our key contributions below:
+  • We introduce FMFA to explicitly leverage fine-grained interactions for improving cross-modal alignment, without incurring extra supervision or inference overhead.
+  • We present an adaptive similarity distribution matching module aimed at precisely aligning image-text pairs in a shared embedding space.
 It adaptively adjusts to narrow the distance between mismatched positive pairs, ensuring more precise matching.
@@
 2. Related Work
-Text-to-Image Person Retrieval (TIPR) was first introduced by Li et al.
-(Li et al., 2017), who constructed the CUHK-PEDES dataset.
-The core challenge of TIPR is to learn a common latent space where visual and textual representations are well aligned.
-Existing methods can be broadly categorized into global and local matching approaches.
-Early global methods (Shu et al., 2022; Zheng et al., 2020; Zhu et al., 2021) directly aligned the global representations of images and text in a joint embedding space.
+Text-to-Image Person Retrieval (TIPR) was initially proposed by Li et al.
+(li2017person), who created the CUHK-PEDES dataset.
+Unlike visual-based person retrieval (cheng2022hybrid; cheng2025semantic; he2025exploring; yang2025feature), the core challenge of TIPR lies in constructing a shared latent space that enables coherent alignment between visual and textual representations.
+Existing methods can be typically classified into global and local matching approaches.
+Early global methods (zheng2020dual; zhu2021dssl) directly aligned the global representations of images and text in a joint embedding space.
 Schroff et al.
-(Schroff et al., 2015) proposed a triplet ranking loss to enforce a margin constraint between positive and negative pairs, and Zhang et al.
-(Zhang and Lu, 2018) introduced the CMPM/C loss to minimize the discrepancy between the scalar projection of image-text pairs and their labels.
+(schroff2015facenet) proposed a triplet ranking loss to enforce a margin constraint between positive and negative pairs, and Zhang et al.
+(zhang2018deep) introduced the CMPM/C loss to minimize the discrepancy between the scalar projection of image-text pairs and their labels.
 However, these global methods lack cross-modal fine-grained interactions, which restrict their ability to capture detailed semantic correspondences.
-To address this limitation, early local matching methods (Wang et al., 2020; Gao et al., 2021; Shao et al., 2022) explicitly aligned local visual and textual features to achieve fine-grained cross-modal interactions.
-Nevertheless, they rely on unimodal pre-trained models (e.g., BERT (Devlin et al., 2019) and ResNet (He et al., 2016)), failing to exploit the strong cross-modal alignment capability of recent pre-trained Vision-Language Models (VLMs) (Li et al., 2022; Yan et al., 2023a; Li et al., 2021).
-Recent local matching methods (Ergasti et al., 2025; Lu et al., 2025; Qin et al., 2024; Wu et al., 2024; Huang et al., 2025) have benefited greatly from VLMs and introduced VLMs to enhance cross-modal alignment.
+To address this limitation, early local matching methods (wang2020vitaa; gao2021contextual; shao2022learning) explicitly aligned local visual and textual features to achieve fine-grained cross-modal interactions.
+Nevertheless, they rely on unimodal pre-trained models (e.g., BERT (devlin2019bert) and ResNet (he2016deep)), failing to exploit the strong cross-modal alignment capability of recent pre-trained Vision-Language Models (VLMs) (li22blip; yan2023clip; li2021align).
+Recent local matching methods (ergasti2025mars; lu2025prompt; qin2024noisy; wu2024text; huang2025cross) have benefited greatly from VLMs and introduced VLMs to enhance cross-modal alignment.
 Park et al.
-(Park et al., 2024) utilized a modified Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021) model as the feature extractor and designed a slot attention-based (Locatello et al., 2020) part discovery module to identify discriminative human parts without extra supervision, while Bai et al.
-(Bai et al., 2023b) used the align-before-fuse model (Li et al., 2021) as the backbone and introduced a cross-modal encoder for fine-grained alignment.
+(park2024plot) utilized a modified Contrastive Language-Image Pre-training (CLIP) (radford2021learning) model as the feature extractor and designed a slot attention-based (locatello2020object) part discovery module to identify discriminative human parts without extra supervision, while Bai et al.
+(bai2023rasa) used the align-before-fuse model (li2021align) as the backbone and introduced a cross-modal encoder for fine-grained alignment.
 Although effective, these methods involve complex computations during inference, leading to high time and memory costs, which limit their applicability to real-time systems.
-On another line of research, several studies (Yang et al., 2023; Shao et al., 2023; Tan et al., 2024) have explored leveraging large-scale image-text pairs in the person Re-IDentification (ReID) domain to VLMs.
+On another line of research, several studies (yang2023towards; shao2023unified; tan2024harnessing) have explored leveraging large-scale image-text pairs in the person Re-IDentification (ReID) domain to VLMs.
 Zuo et al.
-(Zuo et al., 2024) utilized CUHK-PEDES and ICFG-PEDES to train an image captioner, aiming to generate comprehensive textual descriptions for pedestrian images.
+(zuo2024plip) utilized CUHK-PEDES and ICFG-PEDES to train an image captioner, aiming to generate comprehensive textual descriptions for pedestrian images.
 Yang et al.
-(Yang et al., 2023) employed BLIP-2 (Li et al., 2023) to produce attribute-aware captions for diffusion-generated pedestrian images (Rombach et al., 2022), while Jiang et al.
-(Jiang et al., 2025) leveraged recent multi-modal large language models (MLLMs), such as Qwen-VL (Bai et al., 2023a) and LLaVA (Liu et al., 2024), to automatically annotate large-scale ReID datasets in a human-like manner.
+(yang2023towards) employed BLIP-2 (li2023blip2) to produce attribute-aware captions for diffusion-generated pedestrian images (rombach2022high), while Jiang et al.
+(jiang2025modeling) leveraged recent Multi-modal Large Language Models (MLLMs), such as Qwen-VL (bai2023qwenvl) and LLaVA (liu2024llavanext), to automatically annotate large-scale ReID datasets in a human-like manner.
 The CLIP models pre-trained on large-scale ReID datasets exhibit strong zero-shot performance.
 Their compatibility with global matching methods—which relies solely on global features and has a simple inference pipeline—makes them particularly suitable for direct fine-tuning in such settings.
-Among the global matching methods, IRRA (Jiang and Ye, 2023) recently attempted to enhance fine-grained interactions by introducing an Implicit Relation Reasoning (IRR) module, utilizing attention mechanisms to capture cross-modal relations.
-However, this implicit alignment cannot guarantee that all corresponding local features are correctly matched.
+Recent global matching methods (shu2022see; he23vgsg; jiang2023cross) have integrated local fine-grained alignment modules into global matching frameworks to obtain more discriminative global representations.
+Shu et al.
+(shu2022see) introduced a bidirectional mask modeling mechanism that randomly masks image patches and text words, encouraging the model to infer missing semantics and implicitly learn local visual-textual correspondences.
+He et al.
+(he23vgsg) proposed the Vision-Guided Semantic-Group (VGSG) network to cluster textual tokens into semantic groups and align them with corresponding visual regions under the guidance of vision features, achieving group-level fine-grained alignment within a global representation space.
+Similarly, Jiang et al.
+(jiang2023cross) developed IRRA to employ an Implicit Relation Reasoning (IRR) module based on attention mechanisms to capture latent cross-modal relations, enhancing global alignment.
+Although these methods enhance fine-grained cross-modal interactions within global matching frameworks, their implicit or group-level alignment strategies may still fail to guarantee precise local correspondences.
 In light of these limitations, we propose FMFA, which aims to enhance the global matching ability of the model by achieving cross-modal full-mode fine-grained alignment, including explicit fine-grained alignment and implicit relation reasoning.
@@
 Figure 2.
-Overview of the proposed FMFA framework.
-(a) Our FMFA consists of a dual-stream feature extraction backbone and four representation learning branches, i.e.
-Identity Identification (ID loss), Adaptive Similarity Distribution Matching (A-SDM), Explicit Fine-grained Alignment (EFA) and Implicit Relation Reasoning (IRR).
+The illustration of our FMFA framework.
+(a) Our FMFA contains a two-stream feature extraction network and four distinct modules for structured representation learning, namely Identity Identification (ID loss), Adaptive Similarity Distribution Matching (A-SDM), Explicit Fine-grained Alignment (EFA) and Implicit Relation Reasoning (IRR).
 The former two are coarse-grained alignment modules and the latter two are fine-grained alignment modules.
-Modules connected by dashed lines will be removed during inference stage.
+Modules linked via dashed connections are omitted during inference.
 (b) A-SDM obtains the adaptive weight through the similarity score of unmatched positive pairs, adaptively pulling positive pairs closer and dynamically adjusting the pulling force.
@@
 3. Method
-In this section, we present our proposed FMFA framework. The overview of FMFA is illustrated in Figure 2 and the details are introduced in the following subsections.
+This section introduces the proposed FMFA framework. Figure 2 presents an overview of FMFA, and further details of the framework are elaborated in the subsequent subsections.
 3.1. Feature Extraction
-Inspired by the success of IRRA (Jiang and Ye, 2023), we use the modified full CLIP (Radford et al., 2021) visual and textual encoders to enhance cross-modal alignment capabilities while reducing inference costs.
+Motivated by the success of IRRA (jiang2023cross), we use the modified full CLIP (radford2021learning) visual and textual encoders to enhance cross-modal alignment capabilities while reducing inference costs.
 Visual Modality.
-Given an input image $I\in R^{H\times W\times C}$, a CLIP pre-trained ViT model is adopted to obtain the image embeddings.
-We first split $I$ into a sequence of $N=H\times W/P^{2}$ fixed-sized non-overlapping patches, where $P$ denotes the patch size, and then map the patch sequences to 1D tokens $\left\{f_{i}^{v}\right\}|_{i=1}^{N}$ by a trainable linear projection.
-With the injection of the positional embedding and the extra [CLS] token, the sequence of tokens $\left\{f_{cls}^{v},f_{1}^{v},...,f_{N}^{v}\right\}$ are input into $L$-layer transformer blocks to model the correlations of each patch.
-Finally, a linear projection is adopted to map $f_{cls}^{v}$ to the joint image-text embedding space, which serves as the global image representation.
+Given an input image $I\in\mathbb{R}^{H\times W\times C}$, we employ a CLIP-pretrained Vision Transformer (ViT) to attain its image representation.
+An image is first divided into $N=H\times W/P^{2}$ distinct patches of size $P\times P$, which are then transformed into one-dimensional token embeddings $\{f_{i}^{v}\}_{i=1}^{N}$ via a learnable linear projection.
+After adding positional encodings and a [CLS] token, the sequence $\{f_{cls}^{v},f_{1}^{v},\dots,f_{N}^{v}\}$ is passed through $L$ transformer layers to capture dependencies among patches.
+Finally, the [CLS] token embedding $f_{cls}^{v}$ is linearly mapped into the joint image-text embedding space, producing the compact global feature of the image.
 Textual Modality.
-For an input text $T$, we directly use the CLIP-Xformer textual encoder (Radford et al., 2021) to extract the text embeddings.
-Following CLIP, the lower-cased byte pair encoding (Sennrich et al., 2016) is firstly employed to tokenize the input text description.
-The text description is enclosed with [SOS] and [EOS] tokens to indicate the start and end of the sequence.
-The tokenized text $\{f_{sos}^{t},f_{1}^{t},...,f_{eos}^{t}\}$ is then fed into the textual encoder and exploits the correlations of each patch by masked self-attention.
-The output of the highest layer of the transformer at the [EOS] token, $f_{eos}^{t}$, is linearly projected into the image-text joint embedding space to obtain the global text representation.
+Given an input text $T$, we utilize the CLIP-Xformer textual extractor (radford2021learning) to obtain its embedding.
+The text is first tokenized through lower-cased Byte Pair Encoding (BPE) (sennrich2015neural) and framed with [SOS] and [EOS] tokens to indicate sequence boundaries.
+The resulting token sequence $\{f_{sos}^{t},f_{1}^{t},\dots,f_{eos}^{t}\}$ is processed by the transformer encoder, which models dependencies among tokens via masked self-attention.
+Finally, the [EOS] token embedding from the top layer, $f_{eos}^{t}$, is linearly mapped into the joint image-text representation space, generating a compressed global textual representation.
 3.2. Adaptive Similarity Distribution Matching
-Adopted from IRRA (Jiang and Ye, 2023), we introduce a novel Adaptive Similarity Distribution Matching (A-SDM) module, which aims to adaptively pull the unmatched positive image-text pairs into a joint embedding space, thereby enhancing the cross-modal global matching capability of model.
-Given a mini-batch of $B$ image-text pairs, for each global text representation $g_{i}^{t}$ and global image representation $g_{j}^{v}$, we construct a set of image-text representation pairs as $\{(g_{i}^{t},g_{j}^{v}),y_{i,j}\}_{j=1}^{N}$, where $y_{i,j}$ is the matching label.
-$y_{i,j}=1$ means that $(g_{i}^{t},g_{j}^{v})$ is a matched pair, while $y_{i,j}=0$ means that $(g_{i}^{t},g_{j}^{v})$ is an unmatched pair.
-Let $cos(\mathbf{u},\mathbf{v})=\mathbf{u}^{\top}\mathbf{v}/\|\mathbf{u}\|\|\mathbf{v}\|$ denotes the cosine similarity of $u$ and $v$.
-Then, the similarity matrix of matching pairs can be simply calculated by the following softmax function:
-(1) $p_{i,j}=\frac{exp(cos(g_{i}^{t},g_{j}^{v})/\tau_{1})}{\sum_{k=1}^{N}exp(cos(g_{i}^{t},g_{k}^{v})/\tau_{1})},$
-where $\tau_{1}$ is a temperature hyperparameter which controls the probability distribution peaks.
-The matching probability $p_{i,j}$ can be viewed as the proportion of the cosine similarity score between $g_{i}^{t}$ and $g_{j}^{v}$ to the sum of cosine similarity scores between $g_{i}^{t}$ and $\{g_{j}^{v}\}_{j=1}^{N}$ in a mini-batch.
-Let the ithi-th text $T_{i}$ in the mini-batch be the query text and $I_{i}$ be the corresponding image for $T_{i}$ at rank-$k$, where $k>1$.
-Different from IRRA (Jiang and Ye, 2023), we propose to calculate adaptive weight based on the similarity between the query text $T_{i}$ and all image representations:
+Adopted from IRRA (jiang2023cross), we introduce a novel Adaptive Similarity Distribution Matching (A-SDM) module, which aims to adaptively pull the unmatched positive image-text pairs into a shared representation space, further enhancing the cross-modal global matching capability of the model.
+Let the mini-batch contain $B$ image-text pairs, we pair each text embedding $g_{i}^{t}$ with its global image embedding $g_{j}^{v}$ to form the set $\{(g_{i}^{t},g_{j}^{v}),y_{i,j}\}_{j=1}^{B}$, where $y_{i,j}$ serves as the matching indicator.
+Specifically, $y_{i,j}=1$ denotes a matched pair, while $y_{i,j}=0$ denotes an mismatched pair.
+Let $cos(\mathbf{a},\mathbf{c})=\mathbf{a}^{\top}\mathbf{c}/\|\mathbf{a}\|\|\mathbf{c}\|$ denotes the similarity of $\mathbf{a}$ and $\mathbf{c}$.
+Subsequently, like SDM (jiang2023cross), the similarity matrix of image-text pairs is obtained through the following softmax function:
+(1) $p_{i,j}=\frac{exp(cos(g_{i}^{t},g_{j}^{v})/\tau_{1})}{\sum_{k=1}^{B}exp(cos(g_{i}^{t},g_{k}^{v})/\tau_{1})},$
+where $\tau_{1}$ acts as a temperature term that modulates the spread of the resulting distribution.
+The probability $p_{i,j}$ quantifies how much the similarity between the text embedding $g_{i}^{t}$ and the image embedding $g_{j}^{v}$ contributes relative to the sum of all similarities between $g_{i}^{t}$ and every image embedding in the mini-batch.
+Let the ithi-th text $T_{i}$ from the batch be designated as the query text and $I_{i}$ be the corresponding image for $T_{i}$ at rank-$k$, where $k>1$.
+Different from IRRA (jiang2023cross), we propose to derive an adaptive weighting factor by assessing the similarity between the query text $T_{i}$ and all image representations:
 (2) $w_{i}^{t2i}=\alpha\cdot\left[\max_{k}p_{i,k}-p_{i,i}\right]+1,$
-where $\alpha$ is a weight factor which reflects the contribution of unmatched image-text pairs to the cross-modal global matching capability of the model.
-$w_{i}^{t2i}=1$ indicates that the query text $T_{i}$ and its corresponding image $I_{i}$ are successfully matched, while $w_{i}^{t2i}>1$ indicates that $T_{i}$ and $I_{i}$ are unmatched.
-The A-SDM loss from text to image within a mini-batch is then computed as:
+where $\alpha$ is a weight factor reflecting the contribution of unmatched image-text pairs to the cross-modal global matching ability of the model.
+Here, $\max_{k}p_{i,k}$ indicates the top similarity value between the text $T_{i}$ and every image within the mini-batch, while $p_{i,i}$ refers to the similarity associated with its corresponding positive image.
+The constant term “+1+1” ensures that when $T_{i}$ and its corresponding image $I_{i}$ are correctly matched, the weight $w_{i}^{t2i}$ defaults to 1.
+In this case, the A-SDM loss reduces to the SDM loss (jiang2023cross), preventing overemphasis on correctly matched pairs while allowing the model to focus adaptively on harder and misaligned pairs.
+Conversely, $w_{i}^{t2i}>1$ indicates that $T_{i}$ and $I_{i}$ are unmatched, increasing their contribution to the loss to enhance global cross-modal alignment.
+The A-SDM loss for mapping text to image within a mini-batch is subsequently formulated as:
 (3) $\mathcal{L}_{t2i}=W^{t2i}*KL(\mathbf{p_{i}\|q_{i}})\\ =\frac{1}{B}\sum_{i=1}^{B}w_{i}^{t2i}\sum_{j=1}^{B}p_{i,j}\log(\frac{p_{i,j}}{q_{i,j}+\epsilon}),$
-where $\epsilon$ is a small constant to avoid numerical problems, and $q_{i,j}=y_{i,j}/\sum_{k=1}^{B}$ is the true matching probability.
-Symmetrically, the A-SDM loss from image to text $\mathcal{L}_{i2t}$ can be obtained by exchanging the visual and textual modalities. The bi-directional A-SDM loss is calculated by:
+where $\epsilon$ is a tiny offset added to safeguard the computation from unstable values, and $q_{i,j}=y_{i,j}/\sum_{k=1}^{B}$ denotes the ground-truth matching probability.
+In a complementary manner, the A-SDM loss for the image-to-text branch $\mathcal{L}_{i2t}$ is derived by swapping the roles of the text and image features.
+The bi-directional A-SDM loss is formulated as:
 (4) $\mathcal{L}_{A-sdm}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}.$
 3.3. Explicit Fine-grained Alignment
-To fully exploit fine-grained information, it is crucial to bridge the inherent modality differences between visual and textual modalities.
-Although most existing fine-grained alignment methods based on attention mechanisms have achieved certain results by implicitly aligning local features of images and texts, they cannot directly observe whether the corresponding local features of image and text are effectively aligned.
-We introduce an explicit cross-modal aggregation method based on the sparse similarity matrix between the image and text local features.
+To effectively leverage fine-grained information, it is necessary to narrow the underlying disparity between visual and textual modalities.
+Although many attention-based fine-grained alignment approaches have shown effectiveness by implicitly associating local regions in images with textual fragments, they provide no direct means to verify whether these localized correspondences are accurately aligned.
+We propose an explicit cross-modal aggregation approach that leverages the sparse similarity matrix between the local image and text features.
 To further reduce redundant information and minimize memory and time costs during fine-grained alignment, we use hard coding to align the aggregated language-grouped vision embeddings with both image and text embeddings, as shown in Figure 2(c).
+(a) Sparse similarity matrix aggregation.
+(b) Hard coding alignment.
+Figure 3.
+Illustration of the EFA module.
+(a) EFA imposes a sparse structure on the similarity matrix relating token embeddings to patch embeddings, and obtains the language-grouped vision embeddings by aggregating the sparse similarity with its corresponding patch.
+(b) EFA obtains hard similarity through hard coding and LSE pooling, where the calculated hard similarity reflects the relationship between the language-grouped vision embeddings and their original token or patch embeddings.
 Sparse Similarity Matrix Aggregation.
-Some methods (Mukhoti et al., 2023; Yao et al., 2022) use loss functions based on fine-grained representations for cross-modal local alignment, which compute the similarity between all image patch embeddings and text token embeddings.
-Obviously, this approach is computationally and memory-intensive, making it difficult to scale to large batches.
-Therefore, some methods attempt to sparsify the full-size similarity matrix and only consider the retrieved embeddings that are similar to the query embedding.
-Most sparse processing methods are based on softmax, but the softmax approach tends to produce a low-entropy distribution from the perspective of gradient flow (Hoffmann et al., 2024).
-Therefore, we implement sparse similarity matrix aggregation using max-min normalization.
-An image $I$ and its corresponding text $T$ are fed into the visual encoder and textual encoder, respectively, as described in Section 3.1.
-As illustrated in Figure 3(a), the similarity between all image patches and text tokens is obtained by computing the inner product of the last hidden states $\{f_{i}^{t}\}_{i=1}^{L}$ and $\{f_{i}^{v}\}_{i=1}^{N}$ of the text transformer and the vision transformer.
-$s_{i,j}=f_{i}^{t}\cdot f_{j}^{v}$ represents the similarity between the text token $f_{i}^{t}$ and the image patch $f_{j}^{v}$, where $\cdot$ denotes the inner product.
-To obtain the aggregation weight, for each token $i$, we first normalize $s_{i,j}$ to [0,1] using the following min-max normalization formula:
+Some methods (mukhoti2023open; yao2021filip) incur substantial computational and memory overhead, as they evaluate pairwise relationships between every image patch and every text token, which limits scalability to large batch sizes.
+Therefore, we apply a sparsification strategy to reduce the full pairwise similarity computation.
+While softmax is commonly used for such sparse processing, it tends to produce low-entropy similarity distributions that impede effective gradient flow (hoffmann2023eureka).
+Thus, we further adopt a max-min normalization scheme to achieve a more stable and expressive sparse similarity aggregation.
+An image $I$ and its corresponding text $T$ are encoded through the visual and textual encoders, respectively.
+As presented in Figure 3(a), the similarity between image patches and text tokens is computed via the inner product of the last hidden states $\{f_{i}^{t}\}_{i=1}^{L}$ of the text transformer and $\{f_{i}^{v}\}_{i=1}^{N}$ of the vision transformer.
+$s_{i,j}=f_{i}^{t}\cdot f_{j}^{v}$ measures the similarity between the text token $f_{i}^{t}$ and the image patch $f_{j}^{v}$, where $\cdot$ denotes the inner product.
+To obtain the aggregation weight, each token $i$ is first scaled to the range [0,1] through the following min-max normalization:
 (5) $\hat{s}_{i,j}=\frac{s_{i,j}-\min_{k}s_{i,k}}{\max_{k}s_{i,k}-\min_{k}s_{i,k}}.$
@@
 where $\sigma$ is the sparsity threshold.
-$\sigma$ is set to $1/N$, where $N$ is the number of image patches.
+$\sigma$ is assigned the value $1/N$, where $N$ corresponds to the total count of patches in the image.
 This ensures that each token has a minimum of one corresponding image patch for alignment.
@@
 (7) $agg_{i,j}=\frac{\tilde{s}_{i,j}}{\sum_{m=1}^{M}\tilde{s}_{i,j}},$
-where $M$ is the number of image patches retained with high similarity to the token $i$, and $agg_{i,j}$ represents the weight of patch $j$ for aggregating the language-grouped vision embedding corresponding to token $i$.
+where $M$ is the number of image patches retained with high similarity to the token $i$, and $agg_{i,j}$ quantifies the influence of patch $j$ in forming the language-grouped vision embedding (referred to as joint embedding) associated with token $i$.
 This explicit aggregation strategy ensures a comprehensive interaction between token $i$ and its corresponding patch $j$ during local alignment.
 In particular, the aggregation weight $agg_{i,j}$ effectively captures the semantic relevance between token $i$ and patch $j$, thereby facilitating precise alignment.
-Then, we obtain the corresponding language-grouped vision embedding $e_{i}$ as follows:
+Next, we derive the corresponding joint embedding $e_{i}$ as:
 (8) ${e}_{i}=\sum_{j=1}^{N}\\ agg_{i,j}\cdot{f}_{j}^{v},$
-where $N$ denotes the number of image patches. The resulting set of language-grouped vision embedding ${e}_{i}$ has the same length $L$ as the text token $f_{i}^{t}$.
+where $N$ is the count of image patches. The resulting set of joint embedding ${e}_{i}$ has the same length $L$ as the text token $f_{i}^{t}$.
 Hard Coding Alignment.
-We calculate the similarity between the language-grouped vision embeddings (referred to as joint embeddings) $\{e_{i}\}_{i=1}^{L}$ and their corresponding original text embeddings $\{f_{i}^{t}\}_{i=1}^{L}$ as well as image embeddings $\{f_{i}^{v}\}_{i=1}^{N}$, respectively.
-To reduce both computational and memory costs, we adopt a hard coding similarity computation between the joint embeddings and their corresponding text and image embeddings.
+We calculate the similarity between the joint embeddings $\{e_{i}\}_{i=1}^{L}$ and their corresponding original text embeddings $\{f_{i}^{t}\}_{i=1}^{L}$ as well as image embeddings $\{f_{i}^{v}\}_{i=1}^{N}$, respectively.
+To reduce both computational and memory costs, we adopt a hard coding similarity computation between the joint embeddings and their corresponding text and image embeddings, and the theoretical analysis of the hard coding is provided in Appendix A.
 For simplicity, we only present the calculation between the joint embeddings and the text embeddings, while the remaining computations follow a similar and symmetric approach.
-(a) Sparse similarity matrix aggregation.
-(b) Hard coding alignment.
-Figure 3.
-Illustration of the EFA module.
-(a) EFA sparifies the similarity matrix between token embeddings and patch embeddings, and obtains the language-grouped vision embeddings by aggregating the sparse similarity with its corresponding patch.
-(b) EFA obtains hard similarity through hard coding and LSE pooling, where the calculated hard similarity reflects the relationship between the language-grouped vision embeddings and their original token or patch embeddings.
 For the text $T$ and its corresponding joint embedding $E$, we calculate the original similarity matrix $O$ between all text tokens $\{f_{i}^{t}\}_{i=1}^{L}$ and their joint embeddings $\{e_{i}\}_{i=1}^{L}$, where $o_{i,j}={f_{i}^{t}}{e_{j}}^{\top}/\|f_{i}^{t}\|\|e_{j}\|$ means the cosine similarity of $f_{i}^{t}$ and $e_{j}$.
@@
 (9) $\omega_{i,j}=\begin{cases}1&\mathrm{~if~}j=\underset{j^{\prime}=1\cdots L}{\operatorname*{argmax}}(o_{i,j^{\prime}})\\ 0&otherwise\end{cases}.$
-Then, we utilize the LSE pooling (Lee et al., 2018) to compute the hard similarity between text $T$ and its corresponding joint embedding $E$ by:
+Then, we utilize the LSE pooling (lee2018stacked) to compute the hard similarity between text $T$ and its corresponding joint embedding $E$ by:
 (10) $\begin{split}hard\_s(T,E)&=LSE-Pooling\left(\sum_{j=1}^{L}\omega_{i,j}o_{i,j}\right)\\ &=\frac{1}{\lambda}\log\sum_{i=1}^{L}\exp\left(\lambda\max_{j=1\cdots L}o_{i,j}\right),\end{split}$
-where $\lambda$ is a factor that determines how much to magnify the importance of the most relevant pairs of text embedding and its corresponding joint embedding.
-Given a mini-batch of $B$ texts and its corresponding joint embeddings, we obtain the hard coding similarity matrix Hard_SHard\_S according to Eq.
+where $\lambda$ controls the degree to which the most relevant text embeddings and their corresponding joint embeddings are emphasized.
+Given a batch containing $B$ text embeddings along with their associated joint embeddings, we compute the hard coding similarity matrix Hard_SHard\_S following Eq.
 (9) and Eq.
 (10), as illustrated in Figure 3(b).
-We calculate the EFA loss from the text to its joint embedding, adapted from the triplet ranking loss (Schroff et al., 2015):
+We calculate the EFA loss from the text to its joint embedding, adapted from the triplet ranking loss (schroff2015facenet):
 (11) $\mathcal{L}_{t2e}=\frac{1}{B}\log\sum_{\mathrm{neg}}\exp\left(\frac{Hard\_S_{\mathrm{neg}}-Hard\_S_{\mathrm{pos}}+margin}{\tau_{2}}\right),$
-where $\tau_{2}$ is a temperature hyperparameter that controls the distribution of the loss, and marginmargin is a distance hyperparameter representing the minimum distance difference between positive and negative samples.
+where $\tau_{2}$ is a scaling factor adjusting the spread of the loss, and marginmargin is a distance hyperparameter defining the minimal gap separating positive and negative pairs.
 Similarly, the EFA loss from the joint embedding to its original text can be computed following Eq.
-(11), and we can calculate the EFA loss between image and its joint embedding according to Eq.
+(11), and we can calculate the EFA loss between image and its joint embedding through Eq.
 (9), Eq.
@@
 (11).
-Then, we obtain the full EFA loss by:
+Then, we obtain a full EFA loss by:
 (12) $\mathcal{L}_{efa}=\mathcal{L}_{t2e}+\mathcal{L}_{e2t}+\mathcal{L}_{i2e}+\mathcal{L}_{e2i}.$
-Theoretical analysis of hard coding: Given visual embeddings of an image $\boldsymbol{V}\in\mathbb{R}^{N\times d}$ and its joint embeddings $\boldsymbol{E}\in\mathbb{R}^{L\times d}$, where $N$ is the number of image patches and $L$ is the length of the joint embeddings, both the hard coding method and the soft coding method calculate the similarity matrix $A$ to obtain the final similarity between $\boldsymbol{V}$ and $\boldsymbol{E}$, which requires the same time complexity $\mathcal{O}(NLd)$.
-When processing the similarity matrix $A$, our hard coding method calculates the maximum similarity for each row, which has a time complexity of $\mathcal{O}(N)$, as described in Eq.
-(9) and Eq.
-(10).
-However, the soft coding method processes all similarities, resulting in a time complexity of $\mathcal{O}(NL)$.
-Moreover, the spatial complexity of the hard coding method is $\mathcal{O}(NL)$, which is significantly lower than the spatial complexity of the soft coding method, $\mathcal{O}(Ld)$.
-This is based on the fact that $N\ll d$.
-Therefore, our hard coding method can effectively reduce the cost of training time and memory.
 3.4. Training Objective
-As mentioned, the primary goal of FMFA is to enhance the global and cross-modal local alignment of image-text representations within the joint embedding space.
-To achieve it, the most commonly used ID loss (Zheng et al., 2020) and IRR loss (Jiang and Ye, 2023), together with the proposed EFA loss and the A-SDM loss, are jointly employed to optimize FMFA.
-The ID loss directly classifies the global features extracted from each image and text according to their identities, thereby enhancing the global alignment of the model.
-The IRR loss, based on the Masked Language Modeling (MLM) task (Taylor, 1953), leverages an attention mechanism for implicit cross-modal interaction to obtain a joint embedding, and then predicts the [MASK] text token to enhance the local alignment of the model.
-FMFA is trained in an end-to-end manner and the overall training objective is defined as:
+As mentioned, FMFA aims to improve both the global and local cross-modal alignment of image-text features within the shared embedding space.
+To realize this goal, the widely adopted ID loss (zheng2020dual) and IRR loss (jiang2023cross), together with the proposed EFA and A-SDM loss, are jointly utilized to train FMFA.
+The ID loss directly classifies the global features obtained from both the image and the text according to their identities, thereby enhancing the global alignment of the model.
+The IRR loss, based on the Masked Language Modeling (MLM) task (taylor1953cloze), leverages an attention mechanism for implicit cross-modal interaction to obtain a joint embedding, and then predicts the [MASK] text token to enhance the local alignment of the model.
+FMFA is trained end-to-end, with the complete training objective formulated as:
 (13) $\mathcal{L}=\mathcal{L}_{id}+\mathcal{L}_{irr}+\mathcal{L}_{efa}+\mathcal{L}_{A-sdm}.$
@@
 4.1. Datasets and Settings
-In this section, we present the datasets, evaluation metrics and implementation details in our experiments.
 Datasets.
-We use the CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021), and RSTPReid (Zhu et al., 2021) datasets to train and evaluate our FMFA.
-We follow the data partitions used in IRRA (Jiang and Ye, 2023) to split each dataset into training, validation and test sets, where the ICFG-PEDES dataset only has training and validation sets.
-  • CUHK-PEDES is the first benchmark specifically designed for text-to-image person retrieval.
-It comprises 40,206 images and 80,412 textual descriptions corresponding to 13,003 identities.
-According to the official split, the training set includes 11,003 identities, with 34,054 images and 68,108 descriptions.
-The validation and test sets contain 3,078 and 3,074 images, as well as 6,158 and 6,156 textual descriptions, respectively, each covering 1,000 distinct identities.
-  • ICFG-PEDES consists of 54,522 images corresponding to 4,102 identities, with each image paired with a single textual description.
-The dataset is split into a training set and a test set: the training set includes 34,674 image-text pairs from 3,102 identities, while the test set comprises 19,848 pairs covering the remaining 1,000 identities.
-  • RSTPReid comprises 20,505 images of 4,101 identities captured by 15 different cameras.
-Each identity is associated with 5 images taken from different viewpoints, and each image is annotated with 2 textual descriptions.
-According to the official split, the training, validation, and test sets contain 3,701, 200, and 200 identities, respectively.
+We assess FMFA on three widely used text-based person retrieval datasets, following the data splits introduced in IRRA (jiang2023cross).
+CUHK-PEDES (li2017person) contains 40,206 images associated with 13,003 identities, where each image is paired with two textual descriptions.
+Of these identities, 11,003 are designated for training, while the remaining 1,000 identities are allocated separately to validation and test sets.
+ICFG-PEDES (ding2021semantically) includes 54,522 images belonging to 4,102 individuals, each image linked to a single sentence.
+The conventional setup utilizes 3,102 identities for training and reserves 1,000 identities for testing.
+RSTPReid (zhu2021dssl) comprises 20,505 images from 4,101 identities captured across 15 camera views.
+Every identity corresponds to five images taken from different viewpoints, and each image is annotated with two descriptive captions.
+The dataset follows a split with 3,701 identities for training and 200 identities each for validation and testing.
 Evaluation Metrics.
-We mainly employ the popular Rank-K metrics (K=1, 5, 10) to measure the retrieval performance.
-In addition to Rank-K, we also adopt the mean Average Precision (mAP) as an auxiliary retrieval metric to further evaluate performance.
-Higher evaluation metrics indicate better performance of the model.
+To gauge retrieval quality, we primarily report Rank-K results (K = 1, 5, 10), which measure how often the correct item appears within the top-K predictions.
+Additionally, mean Average Precision (mAP) is adopted to summarize ranking accuracy over all query outcomes.
+In both cases, higher metric values correspond to superior model behavior.
 […]
@@
 V. to E. 0.1 1.0 0.8
+Table 2. The hardware configuration of our experimental environment.
+Hardware Details
+CPU Intel Xeon Gold 6330
+GPU NVIDIA RTX A6000
+RAM 755 GB DDR4
 4.2. Comparison with State-of-the-Art Methods
-In this section, we present a comparison with state-of-the-art methods (e.g., NAM (Tan et al., 2024) and HAM (Jiang et al., 2025)) on three public benchmark datasets.
-Based on the backbone type, we categorize the baselines into two groups, as shown in Table 2, Table 3, and Table 4: those employing VL-Backbones w/o ReID-domain pre-training and with ReID-domain pre-training.
-Furthermore, according to whether local features are utilized during inference, the baselines are further divided into local matching methods and global matching methods (denoted as “L” and “G” in the “Type” column, respectively).
-Note that the baseline model in Table 2, Table 3 and Table 4 is denoted as IRRAR, which represents the performance of our reimplementation of the IRRA model.
-CLIP refers to the CLIP-ViT-B/16 model fine-tuned with the InfoNCE loss (Oord et al., 2018).
+In this subsection, we provide a comparison with current state-of-the-art methods (e.g., NAM (tan2024harnessing) and HAM (jiang2025modeling)) on three public benchmark datasets.
+The methods are grouped into two types according to their underlying network architecture, as listed in Table 3, Table 4, and Table 5: those using VL-Backbones without ReID-domain pre-training and those incorporating ReID-domain pre-training.
+Furthermore, according to whether local features are utilized during inference, the baselines are further classified into local and global matching methods (denoted as “L”” and “G” in the “Type” column, respectively).
+It should be noted that the baseline model presented in Table 3, Table 4, and Table 5 is referred to as IRRAR, which represents the performance of our reimplementation of the IRRA model.
+CLIP means the ViT-B/16 architecture after fine-tuning under the InfoNCE loss (oord2018representation).
 Table 3.
-Performance comparisons with state-of-the-art methods on the CUHK-PEDES dataset.
+Comparisons with state-of-the-art methods on the CUHK-PEDES dataset.
 “G”, “L” and “P” in the “Type” column stand for global-matching method, local-matching method and pre-trained model with ReID-domain respectively.
@@
 VL-Backbones w/o ReID-domain pre-training:
-G LGUR (Shao et al., 2022) MM22 ResNet50 BERT 65.25 83.12 89.00 -
-IVT (Shu et al., 2022) ECCV22 ViT-B/16 BERT 65.59 83.11 89.21 -
-CLIP (Radford et al., 2021) ICML21 CLIP-ViT CLIP-Xformer 68.19 86.47 91.47 61.12
-DM-Adapter (Liu et al., 2025) AAAI25 CLIP-ViT CLIP-Xformer 72.17 88.74 92.85 64.33
-IRRAR (Jiang and Ye, 2023) CVPR23 CLIP-ViT CLIP-Xformer 73.45 89.38 93.69 66.13
-TBPS-CLIP (Cao et al., 2024) AAAI24 CLIP-ViT CLIP-Xformer 73.54 88.19 92.35 65.38
+G LGUR (shao2022learning) MM22 ResNet50 BERT 65.25 83.12 89.00 -
+IVT (shu2022see) ECCV22 ViT-B/16 BERT 65.59 83.11 89.21 -
+VGSG (he23vgsg) TIP23 ResNet50 Transformer 67.52 84.37 90.26 -
+CLIP (radford2021learning) ICML21 CLIP-ViT CLIP-Xformer 68.19 86.47 91.47 61.12
+DM-Adapter (liu2025dm) AAAI25 CLIP-ViT CLIP-Xformer 72.17 88.74 92.85 64.33
+IRRAR (jiang2023cross) CVPR23 CLIP-ViT CLIP-Xformer 73.45 89.38 93.69 66.13
+TBPS-CLIP (cao2024empirical) AAAI24 CLIP-ViT CLIP-Xformer 73.54 88.19 92.35 65.38
 FMFA (ours) CLIP-ViT CLIP-Xformer 74.16 90.12 94.10 66.66
-L ACSA (Ji et al., 2022) TMM22 Swin-B BERT 63.56 81.49 87.70 -
-Han et al. (Han et al., 2021) arXiv21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08
-PLOT (Park et al., 2024) ECCV24 CLIP-ViT CLIP-Xformer 75.28 90.42 94.12 -
-RaSa (Bai et al., 2023b) IJCAI23 CLIP-ViT BERT-base 76.51 90.29 94.25 69.38
-PTMI (Lu et al., 2025) TIFS25 CLIP-ViT CLIP-Xformer 76.02 89.93 94.14 70.85
-APTM (Yang et al., 2023) MM23 Swin-B BERT-base 76.53 90.04 94.15 66.91
-SCVD (Wei et al., 2024) TCSVT24 CLIP-RN50 CLIP-Xformer 76.72 90.38 94.89 -
+L ACSA (ji2022asymmetric) TMM22 Swin-B BERT 63.56 81.49 87.70 -
+Han et al. (han2021text) arXiv21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08
+PLOT (park2024plot) ECCV24 CLIP-ViT CLIP-Xformer 75.28 90.42 94.12 -
+RaSa (bai2023rasa) IJCAI23 CLIP-ViT BERT-base 76.51 90.29 94.25 69.38
+PTMI (lu2025prompt) TIFS25 CLIP-ViT CLIP-Xformer 76.02 89.93 94.14 70.85
+APTM (yang2023towards) MM23 Swin-B BERT-base 76.53 90.04 94.15 66.91
+SCVD (wei2024fine) TCSVT24 CLIP-RN50 CLIP-Xformer 76.72 90.38 94.89 -
 VL-Backbones with ReID-domain pre-training:
-P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) ICCV23 CLIP-ViT CLIP-Xformer 74.37 89.51 93.97 66.60
-PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) NeurIPS24 CLIP-ViT CLIP-Xformer 74.25 89.49 93.68 66.52
-NAM (Tan et al., 2024) + IRRAR CVPR24 CLIP-ViT CLIP-Xformer 76.67 91.11 94.60 68.42
-NAM (Tan et al., 2024) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.23 91.33 94.75 68.53
-HAM (Jiang et al., 2025) + IRRAR CVPR25 CLIP-ViT CLIP-Xformer 77.32 91.20 94.95 68.87
-HAM (Jiang et al., 2025) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.46 91.36 95.01 68.89
-Performance Comparisons on CUHK-PEDES.
-We evaluate FMFA on the widely used CUHK-PEDES benchmark, as shown in Table 2.
-When using the VL-Backbones without ReID-domain pre-training, FMFA outperforms all state-of-the-art global matching methods with 74.16% Rank-1 accuracy and 66.66% mAP, and further exceeds IRRA by 0.74% and 0.41% in Rank-5 and Rank-10, respectively.
+P+G UniPT (shao2023unified) + IRRA (jiang2023cross) ICCV23 CLIP-ViT CLIP-Xformer 74.37 89.51 93.97 66.60
+PLIP (zuo2024plip) + IRRA (jiang2023cross) NeurIPS24 CLIP-ViT CLIP-Xformer 74.25 89.49 93.68 66.52
+NAM (tan2024harnessing) + IRRAR CVPR24 CLIP-ViT CLIP-Xformer 76.67 91.11 94.60 68.42
+NAM (tan2024harnessing) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.23 91.33 94.75 68.53
+HAM (jiang2025modeling) + IRRAR CVPR25 CLIP-ViT CLIP-Xformer 77.32 91.20 94.95 68.87
+HAM (jiang2025modeling) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.46 91.36 95.01 68.89
+Table 4. Comparisons with state-of-the-art methods on the RSTPReid dataset.
+Type Method Rank-1 Rank-5 Rank-10 mAP
+VL-Backbones w/o ReID-domain pre-training:
+G DSSL (zhu2021dssl) 39.05 62.60 73.95 -
+IVT (shu2022see) 46.70 70.00 78.80 -
+CLIP (radford2021learning) 54.05 80.70 88.00 43.41
+IRRAR (jiang2023cross) 59.50 81.80 88.85 47.44
+DM-Adapter (liu2025dm) 60.00 82.10 87.90 47.37
+FMFA (ours) 61.05 83.85 89.80 48.22
+L ACSA (lu2025prompt) 48.40 71.85 81.45 -
+CFine (yan2023clip) 50.55 72.50 81.60 -
+PLOT (park2024plot) 61.80 82.85 89.45 -
+RaSa (bai2023rasa) 66.90 86.50 91.35 52.31
+APTM (yang2023towards) 67.50 85.70 91.45 52.56
+VL-Backbones with ReID-domain pre-training:
+P+G UniPT (shao2023unified) + IRRA (jiang2023cross) 62.20 83.30 89.75 48.33
+PLIP (zuo2024plip) + IRRA (jiang2023cross) 64.35 83.75 91.00 50.93
+NAM (tan2024harnessing) + IRRAR 68.25 86.75 92.30 52.92
+NAM (tan2024harnessing) + FMFA (ours) 68.70 87.05 92.35 53.14
+HAM (jiang2025modeling) + IRRAR 71.35 87.60 93.05 55.40
+HAM (jiang2025modeling) + FMFA (ours) 71.80 88.05 93.15 55.72
+Evaluation Results on CUHK-PEDES We measure the performance of FMFA on the CUHK-PEDES dataset, as presented in Table 3.
+When using the VL-Backbones without ReID-domain pre-training, FMFA achieves superior performance over advanced global matching methods, attaining 74.16% Rank-1 and 66.66% mAP, while surpassing IRRA by 0.74% in Rank-5 and 0.41% in Rank-10.
 When adopting the VL-Backbones with ReID-domain pre-training, FMFA maintains its superiority, and achieves Rank-5 accuracy exceeding 95% with the HAM-based backbone.
 Notably, FMFA with NAM-based backbone attains 91.33% in Rank-5, outperforming IRRA with the HAM-based backbone by 0.13%.
-Table 3. Performance comparisons with state-of-the-art methods on the RSTPReid dataset.
-Type Method Rank-1 Rank-5 Rank-10 mAP
-VL-Backbones w/o ReID-domain pre-training:
-G DSSL (Zhu et al., 2021) 39.05 62.60 73.95 -
-IVT (Shu et al., 2022) 46.70 70.00 78.80 -
-CLIP (Radford et al., 2021) 54.05 80.70 88.00 43.41
-IRRAR (Jiang and Ye, 2023) 59.50 81.80 88.85 47.44
-DM-Adapter (Liu et al., 2025) 60.00 82.10 87.90 47.37
-FMFA (ours) 61.05 83.85 89.80 48.22
-L ACSA (Lu et al., 2025) 48.40 71.85 81.45 -
-CFine (Yan et al., 2023a) 50.55 72.50 81.60 -
-PLOT (Park et al., 2024) 61.80 82.85 89.45 -
-RaSa (Bai et al., 2023b) 66.90 86.50 91.35 52.31
-APTM (Yang et al., 2023) 67.50 85.70 91.45 52.56
-VL-Backbones with ReID-domain pre-training:
-P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 62.20 83.30 89.75 48.33
-PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 64.35 83.75 91.00 50.93
-NAM (Tan et al., 2024) + IRRAR 68.25 86.75 92.30 52.92
-NAM (Tan et al., 2024) + FMFA (ours) 68.70 87.05 92.35 53.14
-HAM (Jiang et al., 2025) + IRRAR 71.35 87.60 93.05 55.40
-HAM (Jiang et al., 2025) + FMFA (ours) 71.80 88.05 93.15 55.72
-Performance Comparisons on RSTPReid.
-We evaluate FMFA on the newest benchmark, RSTPReid, as shown in Table 3.
+Evaluation Results on RSTPReid.
+We assess FMFA on the latest RSTPReid benchmark, as presented in Table 4.
 Using the VL-Backbones without ReID-domain pre-training, FMFA achieves competitive performance, attaining 61.05% Rank-1, 83.85% Rank-5, 89.80% Rank-10, and 48.22% mAP, respectively, outperforming IRRA by 1.55% in Rank-1 and 2.05% in Rank-5.
@@
 Notably, FMFA achieves Rank-5 accuracy higher than 88% with the HAM-based backbone.
-Table 4. Performance comparisons with state-of-the-art methods on the ICFG-PEDES dataset.
+Table 5. Comparisons with state-of-the-art methods on the ICFG-PEDES dataset.
 Type Method Rank-1 Rank-5 Rank-10 mAP
 VL-Backbones w/o ReID-domain pre-training:
-G Dual Path (Zheng et al., 2020) 38.99 59.44 68.41 -
-IVT (Shu et al., 2022) 56.04 73.60 80.22 -
-CLIP (Radford et al., 2021) 56.74 75.72 82.26 31.84
-DM-Adapter (Liu et al., 2025) 62.64 79.53 85.32 36.50
-IRRAR (Jiang and Ye, 2023) 63.48 80.16 85.78 38.20
+G Dual Path (zheng2020dual) 38.99 59.44 68.41 -
+IVT (shu2022see) 56.04 73.60 80.22 -
+CLIP (radford2021learning) 56.74 75.72 82.26 31.84
+VGSG (he23vgsg) 60.34 76.01 82.01 -
+DM-Adapter (liu2025dm) 62.64 79.53 85.32 36.50
+IRRAR (jiang2023cross) 63.48 80.16 85.78 38.20
 FMFA (ours) 64.29 80.48 85.93 39.43
-L SSAN (Ding et al., 2021) 54.23 72.63 79.53 -
-ISANet (Yan et al., 2023b) 57.73 75.42 81.72 -
-CFine (Yan et al., 2023a) 60.83 76.55 82.42 -
-RaSa (Bai et al., 2023b) 65.28 80.40 85.12 41.29
-PLOT (Park et al., 2024) 65.76 81.39 86.73
+L SSAN (ding2021semantically) 54.23 72.63 79.53 -
+ISANet (yan2023image) 57.73 75.42 81.72 -
+CFine (yan2023clip) 60.83 76.55 82.42 -
+RaSa (bai2023rasa) 65.28 80.40 85.12 41.29
+PLOT (park2024plot) 65.76 81.39 86.73
 VL-Backbones with ReID-domain pre-training:
-P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 64.50 80.24 85.74 38.22
-PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 65.79 81.94 87.32 39.43
-NAM (Tan et al., 2024) + IRRAR 66.34 81.94 86.73 40.14
-NAM (Tan et al., 2024) + FMFA (ours) 66.58 81.94 87.04 40.17
-HAM (Jiang et al., 2025) + IRRAR 68.21 83.28 88.04 41.72
-HAM (Jiang et al., 2025) + FMFA (ours) 68.37 83.28 88.10 41.76
-Performance Comparisons on ICFG-PEDES.
-We also evaluate FMFA on another benchmark, ICFG-PEDES, as shown in Table 4.
-Using VL-Backbones without ReID-domain pre-training, FMFA achieves the best performance across all evaluation metrics, attaining 64.29% Rank-1 and 39.43% mAP.
-Compared with IRRA, FMFA shows a notable improvement of 0.81% in Rank-1 and 1.23% in mAP, which is meaningful for practical applications.
+P+G UniPT (shao2023unified) + IRRA (jiang2023cross) 64.50 80.24 85.74 38.22
+PLIP (zuo2024plip) + IRRA (jiang2023cross) 65.79 81.94 87.32 39.43
+NAM (tan2024harnessing) + IRRAR 66.34 81.94 86.73 40.14
+NAM (tan2024harnessing) + FMFA (ours) 66.58 81.94 87.04 40.17
+HAM (jiang2025modeling) + IRRAR 68.21 83.28 88.04 41.72
+HAM (jiang2025modeling) + FMFA (ours) 68.37 83.28 88.10 41.76
+Evaluation Results on ICFG-PEDES.
+We assess FMFA on the ICFG-PEDES benchmark, with the results displayed in Table 5.
+Using VL-Backbones without ReID-domain pre-training, FMFA obtains the leading results across all metrics, attaining 64.29% Rank-1 and 39.43% mAP.
+Compared with IRRA, FMFA shows a notable improvement of 0.81% Rank-1 and 1.23% mAP, which is meaningful for practical applications.
 When adopting VL-Backbones with ReID-domain pre-training, FMFA yields slight gains, outperforming IRRA by 0.24% and 0.16% in Rank-1 with the NAM-based and HAM-based backbones, respectively.
-In summary, our FMFA achieves the best performance across all evaluation metrics on three common public benchmarks.
-To the best of our knowledge, FMFA is the best method for all global matching methods.
-This demonstrates the generalizability and robustness of our proposed method.
-Table 5. Ablation study on each component of FMFA on CUHK-PEDES, ICFG-PEDES and RSTPReid.
-No. Methods Components CUHK-PEDES ICFG-PEDES RSTPReid
-SDM A-SDM EFA Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10 Rank-1 Rank-5 Rank-10
-0 Baseline $\checkmark$ 73.45 89.38 93.69 63.48 80.16 85.78 59.50 81.80 88.85
-1 +A-SDM $\checkmark$ 74.04 89.86 93.89 64.26 80.59 85.90 60.25 81.45 88.70
-2 +EFA $\checkmark$ $\checkmark$ 73.73 89.64 94.04 63.77 80.39 85.86 60.45 82.30 89.25
-3 FMFA $\checkmark$ $\checkmark$ 74.16 90.12 94.10 64.29 80.48 85.93 61.05 83.85 89.80
+In conclusion, FMFA attains the highest performance across all evaluation metrics on the three widely used public benchmarks.
+As far as we are aware, FMFA is the best method for all global matching methods.
+This highlights the ability of our method to generalize well and maintain robustness.
 4.3. Ablation Study
-In this section, we analyze our proposed components in the FMFA framework.
+In this subsection, we examine our proposed components in the FMFA framework.
 For simplicity, we omit the components of $\mathcal{L}_{id}$ and the IRR module that were proposed by IRRA and used in all experiments.
 Only one of SDM and A-SDM can be used at the same time.
-To fully demonstrate the impact of our proposed components in FMFA, we conduct an empirical analysis on three common datasets (i.e., CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021) and RSTPReid (Zhu et al., 2021)).
-The Rank-1, Rank-5 and Rank-10 accuracies(%) are shown in Table 5.
+To thoroughly assess the contribution of our FMFA modules, we undertake an empirical analysis on three widely used datasets.
+Table 6 summarizes the Rank-1/5/10 accuracies (%) together with the mAP (%) performance.
 Effect of The A-SDM Module.
 To evaluate the contribution of the Adaptive Similarity Distribution Matching (A-SDM) module, we perform ablation experiments by replacing the A-SDM module with the SDM module, keeping all hyperparameters unchanged.
-Specifically, as shown in Table 5, replacing A-SDM with SDM leads to a decrease in the Rank-1 accuracy by 0.59%, 0.78%, and 0.75% across the three datasets, as observed in No.
+Specifically, as shown in Table 6, replacing A-SDM with SDM results in a reduction of Rank-1 accuracy by 0.59%, 0.78%, and 0.75% across the three datasets, and also causes a 1.19% drop in mAP on the ICFG-PEDES dataset, as observed in No.
 0 vs.
@@
 1.
-Additionally, all evaluation metrics on the CUHK-PEDES and ICFG-PEDES datasets degrade, further confirming the superiority of A-SDM.
+Additionally, all evaluation metrics on CUHK-PEDES and ICFG-PEDES degrade, further confirming the superiority of A-SDM.
 Moreover, when combined with the EFA module, the advantage of A-SDM becomes even more pronounced.
@@
 No.
-3, replacing the A-SDM module with the SDM module results in 1.26% and 1.00% decrease in Rank-1 and Rank-5 on the CUHK-PEDES dataset, respectively, and a 1.55% drop in the Rank-5 accuracy on the RSTPReid dataset.
-These results collectively validate the consistent and significant contribution of the A-SDM module to overall performance.
+3, replacing the A-SDM module with the SDM module results in 0.43% and 0.48% decrease in Rank-1 and Rank-5 on the CUHK-PEDES dataset, respectively, as well as a 1.55% drop in Rank-5 and a 0.58% decline in mAP on the RSTPReid dataset.
+These results collectively validate the consistent and significant impact of A-SDM to performance.
+Table 6. Ablation analysis of FMFA modules across three public benchmarks.
+No. Methods Components CUHK-PEDES ICFG-PEDES RSTPReid
+SDM A-SDM EFA Rank-1 Rank-5 Rank-10 mAP Rank-1 Rank-5 Rank-10 mAP Rank-1 Rank-5 Rank-10 mAP
+0 Baseline $\checkmark$ 73.45 89.38 93.69 66.13 63.48 80.16 85.78 38.20 59.50 81.80 88.85 47.44
+1 +A-SDM $\checkmark$ 74.04 89.86 93.89 66.45 64.26 80.59 85.90 39.39 60.25 81.45 88.70 47.69
+2 +EFA $\checkmark$ $\checkmark$ 73.73 89.64 94.04 66.40 63.77 80.39 85.86 39.17 60.45 82.30 89.25 47.64
+3 FMFA $\checkmark$ $\checkmark$ 74.16 90.12 94.10 66.66 64.29 80.48 85.93 39.43 61.05 83.85 89.80 48.22
 Effect of The EFA Module.
-To enhance the global matching capability of model, the Explicit Fine-grained Alignment (EFA) module introduces fine-grained cross-modal interaction based on a sparse similarity matrix.
-The effectiveness of the EFA module is demonstrated through the experimental comparisons between No.
+To improve the model’s global matching performance, the Explicit Fine-grained Alignment (EFA) module introduces fine-grained cross-modal interaction based on a sparse similarity matrix.
+The impact of the EFA module is illustrated by comparing the results of No.
 0 vs.
@@
 No.
-3, removing the EFA module from FMFA leads to a performance drop of 0.26% and 0.21% in Rank-5 and Rank-10 on the CUHK-PEDES dataset, and a more significant decline of 2.40% and 1.10% in Rank-5 and Rank-10 on the RSTPReid dataset.
+3, removing the EFA module from FMFA leads to a performance drop of 0.26% and 0.21% in Rank-5 and Rank-10 on the CUHK-PEDES dataset, and a more significant decline of 2.40% and 1.10% in Rank-5 and Rank-10, along with a 0.53% decrease in mAP on the RSTPReid dataset.
 However, EFA causes a 0.11% drop in Rank-5 on ICFG-PEDES, suggesting that its sparse and hard coding strategy, which focuses only on the most relevant patches, may overlook other informative ones and lead to information loss.
@@
 2.
-In this setting, the absence of EFA results in 0.38%, 0.29%, and 0.95% drops in Rank-1 on CUHK-PEDES, ICFG-PEDES and RSTPReid, respectively.
-These comparisons demonstrate the effectiveness of the EFA module.
+In this setting, the absence of EFA results in 0.38%, 0.29%, and 0.95% drops in Rank-1 on CUHK-PEDES, ICFG-PEDES and RSTPReid, and causes a 1.23% drop in mAP on ICFG-PEDES.
+These evaluations highlight the effectiveness of the EFA module.
+4.4. Parameter Study
+We perform a parameter study on the CUHK-PEDES dataset, examining three hyperparameters—the weight factor $\alpha$, the sparsity threshold $\sigma$, and the factor $\lambda$—as well as the contribution weights of the proposed loss functions, $\mathcal{L}_{A\text{-}SDM}$ and $\mathcal{L}_{EFA}$.
+When examining a specific parameter, all other parameters are maintained as specified in Section 4.1.
 (a) The weight factor $\alpha$ in Eq. (2).
@@
 Figure 4. The sensitivity analysis of hyperparameters of FMFA on the CUHK-PEDES dataset.
-4.4. Parameter Study
-We conduct the parameter analysis on the CUHK-PEDES dataset, including three hyperparameters (i.e., the weight factor $\alpha$, the sparsity threshold $\sigma$ and the factor $\lambda$) and the weight of the proposed loss functions (i.e., $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$).
-While we analyze the special parameter, other parameters are kept the same as those described in Section 4.1.
 Hyperparameters Analysis.
 As shown in Figure 4(a), we vary the weight factor $\alpha$ from 1 to 20.
-The results on the CUHK-PEDES dataset indicate that setting $\alpha$ to 10 yields the best performance across all metrics, suggesting that the adaptive pull force on unmatched positive pairs is optimal.
+The results on CUHK-PEDES show that setting $\alpha$ to 10 achieves the highest performance across the evaluated metrics, suggesting that the adaptive pull force on unmatched positive pairs is optimal.
 However, when $\alpha$ is increased to 20, the performance drops significantly because the pull force becomes excessively strong, causing unmatched positive pairs to fail to align properly and instead over-pulling mismatched positives, which leads to false positives.
@@
 As illustrated in Figure 4(b) and Figure 4(c), setting $\sigma$ to $1/N$ and $\lambda$ to 1 yields the best overall performance.
-Loss Function Weights.
-We conduct a series of experiments on the CUHK-PEDES dataset to analyze the effect of the weights of the two proposed loss functions, $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$, varying them from 0.1 to 10.0, as shown in Table 6.
-The results show that setting both weights to 1 yields the best performance.
-Notably, increasing the weight of $\mathcal{L}_{efa}$ to 5.0 or 10.0 results in gradient explosion, making the model untrainable and causing all metrics to drop below 1.
-Table 6. The sensitivity analysis of the weights of $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$ on the CUHK-PEDES dataset.
+When $\sigma$ is set too high, only a few highly relevant patches are retained, leading to the loss of semantic information, whereas an excessively low $\sigma$ preserves most patches and weakens the ability to capture discriminative features.
+According to Eq.
+(10), a large $\lambda$ emphasizes the differences among patch responses and makes the pooling operation less robust to noise, while a small $\lambda$ smooths these responses excessively and reduces feature discrimination.
+Moreover, to evaluate the stability of these hyperparameters, we further set $\sigma$ around $1/N$ (i.e., $0.9/N$ and $1.1/N$) and $\lambda$ around 1 (i.e., 0.9 and 1.1).
+The results show that although slightly adjusting $\sigma$ and $\lambda$ leads to a minor performance drop, our model still obtains high performance, demonstrating the robustness of our method.
+Table 7. The sensitivity analysis of the weight of $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$ on the CUHK-PEDES dataset, where one weight is varied and the other is fixed at 1.0.
 Weight $\mathcal{L}_{A-sdm}$ $\mathcal{L}_{efa}$
@@
 10.0 72.45 88.62 93.24 0.09 0.19 0.32
+Table 8. Ablation study verifying the necessity of the constant term “+1” in A-SDM on the CUHK-PEDES dataset.
+Setting “+1” Term Rank-1 Rank-5 Rank-10
+A-SDM w/o “+1” 26.12 48.36 59.18
+A-SDM w “+1” $\checkmark$ 74.16 90.12 94.10
+Loss Function Weights.
+We perform experiments using the CUHK-PEDES dataset to investigate the influence of the weights of the two proposed loss functions, $\mathcal{L}_{A-sdm}$ and $\mathcal{L}_{efa}$, varying them from 0.1 to 10.0, as shown in Table 7.
+To comprehensively explore the optimal combination of the proposed loss weights, we only vary one weight at a time while keeping the other fixed at 1 to ensure a controlled comparison.
+The results show that setting both weights to 1 yields the best performance.
+Notably, increasing the weight of $\mathcal{L}_{efa}$ to 5.0 or 10.0 results in gradient explosion, making the model untrainable and causing all metrics to drop below 1.
 Figure 5. Visualization of similarity maps before and after the sparse process in the EFA module.
+The “+1” term in A-SDM.
+We carry out experiments on the CUHK-PEDES dataset to validate the necessity of the constant “+1” in A-SDM.
+As defined in Eq.
+(2), the constant term “+1” ensures a positive lower bound for the adaptive weight, preventing it from becoming zero when a pair is correctly matched.
+Without this term, gradients of correctly matched positive pairs diminish, leading to unstable optimization and degraded performance during training.
+Table 8 indicates that omitting the “+1” leads to a notable decline in retrieval accuracy, illustrating that the effectiveness of model is reduced.
+Therefore, the “+1” term is a crucial component that ensures stable optimization and consistent performance of A-SDM.
 4.5. Qualitative Results
@@
 (6)) in the EFA module, as shown in Figure 5.
-Patches with high similarity (greater than 0.75) are preserved, whereas those with low similarity (below 0.5) are suppressed to values under 0.25 and thus omitted during aggregation (left in Figure 5).
-When most patches in an image exhibit relatively high similarity (around 0.75), the sparse process retains only the patches with the highest similarity, while pushing the similarity of the remaining patches below 0.5 (right in Figure 5).
-This ensures that the subsequent aggregation focuses only on the most relevant patches, reducing computational and memory costs while maintaining performance.
+Patches with high similarity (greater than 0.75) are preserved, as indicated by the black frames in the left part of Figure 5.
+In contrast, patches with low similarity (below 0.5) are suppressed to values under 0.25 and thus omitted during aggregation, as shown by the pink frames on the left.
+When most patches in an image exhibit relatively high similarity (around 0.75), the sparse process retains only those with the highest similarity while reducing the similarity of the remaining patches below 0.5, highlighted by the purple frames in the right part of Figure 5.
+This guarantees that the subsequent aggregation emphasizes only the most pertinent patches, decreasing computational and memory overhead while preserving performance.
 Unlike implicit aggregation methods based on attention mechanisms, EFA explicitly aggregates image patches and text tokens, allowing us to observe whether the most relevant patches are effectively aggregated by visualizing the similarity between patches and tokens.
-Inference Time Comparison We compare the inference time between FMFA and recent local matching methods (e.g., PLOT (Park et al., 2024) and RaSa (Bai et al., 2023b)) on the test sets of three datasets, as shown in Table 7.
+Inference Time Comparison.
+We compare the inference time between FMFA and recent matching methods (e.g., PLOT (park2024plot) and RaSa (bai2023rasa)) on the test sets of three datasets, as shown in Table 9.
 As a global matching method, FMFA only computes global features during inference, thus achieving a higher inference speed than local matching methods.
-Moreover, the efficiency advantage of FMFA becomes more pronounced as the size of the test set increases.
+Even when compared with the recent global matching method DM-Adapter (liu2025dm), FMFA still achieves consistently faster inference across all three datasets.
+Moreover, the efficiency advantage of FMFA becomes more pronounced with the growth of the test set.
 For instance, the inference time of FMFA vs.
@@
 91s on ICFG-PEDES.
-Table 7. Comparison of the inference time (s) between FMFA and recent local matching methods.
+These comparisons clearly demonstrate that FMFA achieves superior inference speed compared with recent methods.
+Table 9. Comparison of the inference time (s) between FMFA and recent methods.
 Method RSTPReid CUHK-PEDES ICFG-PEDES
 FMFA (ours) 3 7 50
-PLOT (Park et al., 2024) 5 16 91
-RaSa (Bai et al., 2023b) 388 1168 3871
+DM-Adapter (liu2025dm) 8 16 78
+PLOT (park2024plot) 5 16 91
+APTM (yang2023towards) 10 29 95
+SCVD (wei2024fine) - 191 1369
+RaSa (bai2023rasa) 388 1168 3871
 (a) A man wearing a gray and black shirt, a pair of black pants and a bag over his right shoulder.
@@
 (d) Female holding her phone with long black hair, wearing jogging clothes and tennis shoes carrying bags.
-Figure 6.
-Comparison of top-5 retrieved results on CUHK-PEDES between baseline and FMFA for each text query.
-The image corresponding to query text, along with the matched and mismatched images, are highlighted with black, green and red rectangles, respectively.
+Figure 6. Top-5 retrieval results for each text query on CUHK-PEDES, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.
 Visualization of Top-5 Retrieval Results.
 Figure 6 compares the top-5 retrieval results between the baseline IRRAR and our proposed FMFA on the CUHK-PEDES dataset.
-As shown in Figure 6, FMFA achieves more accurate retrieval results, successfully retrieving images that the baseline fails to match.
+Figure 6 illustrates that FMFA delivers more precise retrieval results, correctly identifying images that the baseline fails to match.
 For query texts where the baseline performs well, FMFA further improves performance by retrieving more relevant pedestrian images (e.g., Figure 6(a) and Figure 6(c)).
@@
 This is because our proposed FMFA focuses on the unmatched positive pairs and adaptively pulls the positive pairs closer.
-More comparisons of the top-5 retrieved results between the baseline and FMFA on the ICFG-PEDES and RSTPReid datasets are provided in Figure 7 and Figure 8.
+More comparisons of the top-5 retrieved results are provided in Appendix B.
+5. Conclusion
+In summary, we propose a cross-modal Full-Mode Fine-grained Alignment (FMFA) framework to learn discriminative global text-image representations through full-mode fine-grained alignment, including explicit fine-grained alignment and existing implicit relational reasoning.
+We design an Adaptive Similarity Distribution Matching (A-SDM) module to concentrate on unmatched positive pairs, adaptively pulling them closer.
+In addition, to achieve cross-modal fine-grained alignment, we introduce an Explicit Fine-grained Alignment (EFA) module, which explicitly aggregates local text and image representations based on the sparse similarity matrix and employs a hard coding method.
+These modules function together to project images and text into a shared embedding space.
+Comprehensive experiments across three datasets confirm the effectiveness and superior performance of our FMFA framework.
+Limitation.
+The fixed threshold in the sparse process only keeps the most relevant patches, which may result in the loss of semantic information and limit the effective aggregation of local features, thereby affecting the overall performance of our model.
+Incorporating adaptive methods that capture complete semantic information (e.g., tree transformer (wang2019tree)) could further enhance our model.
+Acknowledgements.
+This work was supported by the National Natural Science Foundation of China (No.
+62302080), Guangxi Key Research and Development Program (No.
+Guike AB24010112), National Foreign Expert Project of China (No.
+S20240327), Sichuan Science and Technology Program (No.
+2025HJRC0021) and Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No.
+BX202312).
+Appendix A Theoretical Analysis of Hard Coding in EFA
+Given visual embeddings of an image $\boldsymbol{V}\in\mathbb{R}^{N\times d}$ and its joint embeddings $\boldsymbol{E}\in\mathbb{R}^{L\times d}$, where $N$ is the number of image patches and $L$ is the length of the joint embeddings, both the hard coding method and the soft coding method calculate the similarity matrix $A$ to obtain the final similarity between $\boldsymbol{V}$ and $\boldsymbol{E}$, which requires the same time complexity $\mathcal{O}(NLd)$.
+When processing the similarity matrix $A$, our hard coding method calculates the maximum similarity for each row, which has a time complexity of $\mathcal{O}(N)$, as described in Eq.
+(9) and Eq.
+(10).
+However, the soft coding method processes all similarities, resulting in a time complexity of $\mathcal{O}(NL)$.
+Moreover, the spatial complexity of the hard coding method is $\mathcal{O}(NL)$, which is significantly lower than the spatial complexity of the soft coding method, $\mathcal{O}(Ld)$.
+This is based on the fact that $N\ll d$.
+Therefore, our hard coding method can effectively reduce the cost of training time and memory.
+Appendix B More Comparisons of The Top-5 Retrieved Results
+Figure 7 and Figure 8 show the comparisons of the top-5 retrieved results between the baseline and FMFA on the ICFG-PEDES and RSTPReid datasets, respectively.
+These comparisons further highlight the superiority and effectiveness of the proposed FMFA framework.
 (a) A man in his forties with short black hair is wearing a mid-length grey trench coat over a light-colored collared shirt. He is also wearing a regular-fit black pants.
@@
 (d) A middle-aged man with black medium length hair with receding hairline and he is wearing a green hooded jacket with black pants paired with black shoes.
-Figure 7.
-Comparison of top-5 retrieved results on ICFG-PEDES between baseline and FMFA for each text query.
-The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
+Figure 7. Top-5 retrieval results for each text query on ICFG-PEDES, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.
 (a) A person is wearing a short dark jacket, a pair of green overalls and a pair of sneakers with blue edging.
@@
 (d) A man walking on the street is wearing a navy blue jacket,loose jeans and a pair of sneakers. He puts his jacket’s hat on the head.And a bag and some paper in his hand.
-Figure 8.
-Comparison of top-5 retrieved results on RSTPReid between baseline and FMFA for each text query.
-The image corresponding to query text, matched and mismatched images are marked with black, green and red rectangles, respectively.
-5. Conclusion
-In this paper, we propose a cross-modal Full-Mode Fine-grained Alignment (FMFA) framework to learn discriminative global text-image representations through full-mode fine-grained alignment, including explicit fine-grained alignment and existing implicit relational reasoning.
-We design an Adaptive Similarity Distribution Matching (A-SDM) module to concentrate on unmatched positive pairs, adaptively pulling them closer.
-In addition, to achieve cross-modal fine-grained alignment, we introduce an Explicit Fine-grained Alignment (EFA) module, which explicitly aggregates local text and image representations based on the sparse similarity matrix and employs a hard coding method.
-These modules work collaboratively to align images and text into a joint embedding space.
-Extensive experiments on three datasets demonstrate the superiority and effectiveness of our proposed FMFA framework.
-Limitation.
-The fixed threshold in the sparse process only keeps the most relevant patches, which may result in the loss of semantic information and limit the effective aggregation of local features, thereby affecting the overall performance of our model.
-Incorporating adaptive methods that capture complete semantic information (e.g., tree transformer (Wang et al., 2019)) could further enhance our model.
-Acknowledgements.
-This work was supported by the National Natural Science Foundation of China (No.
-62302080), National Foreign Expert Project of China (No.
-S20240327), Sichuan Science and Technology Program (No.
-2025HJRC0021) and Sichuan Province Innovative Talent Funding Project for Postdoctoral Fellows (No.
-BX202312).
+Figure 8. Top-5 retrieval results for each text query on RSTPReid, comparing baseline with FMFA. Target image, correct matches, and mismatches are outlined in black, green, and red.